### v1

# Assessing the Effects of Monetary Shocks on Macroeconomic Stars: A SMUC-IV Framework

Abstract
This paper proposes a structural multivariate unobserved components model with external instrument (SMUC-IV) to investigate the effects of monetary policy shocks on key U.S. macroeconomic “stars”—namely, the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate. A key feature of our approach is the use of an external instrument to identify monetary policy shocks within the multivariate unobserved components modeling framework. We develop an MCMC estimation method to facilitate posterior inference within our proposed SMUC-IV framework. In addition, we propose an marginal likelihood estimator to enable model comparison across alternative specifications. Our empirical analysis shows that contractionary monetary policy shocks have significant negative effects on the macroeconomic stars, highlighting the nonzero long-run effects of transitory monetary policy shocks.
Keywords: macroeconomic stars, monetary policy, unobserved components models, external instrument, precision sampling, marginal likelihood estimation
JEL classification: C11, C32, E52
1 Introduction
Long-run equilibrium levels of key macroeconomic variables—commonly referred to as “stars”—are central to guiding countercyclical policy and assessing long-run performance. In the conventional view, the macroeconomic stars are unaffected by monetary policy shocks, which are thought to have only transitory effects. Recently, a growing body of evidence challenges this view, showing that monetary shocks can have long-run influences on the economy (e.g., Hanson and Stein, , 2015; Diegel and Nautz, , 2021; Jordà et al., , 2024). A less studied empirical question, however, is how monetary policy shocks affect the macroeconomic stars.
The main objective of this paper is to develop a unified econometric framework to address this question. Multivariate unobserved components models have become the standard econometric tool for analyzing macroeconomic stars (e.g., Kuttner, , 1994; Laubach and Williams, , 2003; Chan et al., , 2016; Zaman, , 2025). These models decompose observed macroeconomic series into long-run trends and short-run cycles, identifying the stars with the trend components. The traditional multivariate unobserved components model framework, albeit useful to model the macroeconomic stars, lacks a formal strategy to identify the effects of monetary policy shocks on the stars.
This paper contributes to the existing literature by introducing a structural multivariate unobserved components model to assess the effects of monetary policy shocks on the macroeconomic stars. The main novelty of our approach is the use of an external instrument to identify monetary policy shocks for structural analysis within a multivariate unobserved components model, which we refer to as SMUC-IV. While the use of external instruments to identify macroeconomic shocks in structural vector autoregressions has become increasingly common (Mertens and Ravn, , 2013; Ramey, , 2016; Stock and Watson, , 2018; Jarociński and Karadi, , 2020; Caldara and Herbst, , 2019; Arias et al., , 2021), very few studies have considered their application within multivariate unobserved components models. The SMUC-IV introduced in this paper addresses this gap. Although the SMUC-IV developed in this paper is used to assess the effects of monetary policy shocks on the macroeconomic stars, the methodology readily extends to broader applications in a straightforward manner.
Our proposed SMUC-IV possesses two desirable characteristics. First, it offers a unified framework for the joint estimation of the macroeconomic stars and the assessment of their responses to monetary policy shocks. This ensures internally consistent estimates of the macroeconomic stars, enabling a coherent analysis of the dynamic interactions between the long-run trends and short-run cyclical fluctuations. Second, our SMUC-IV specification allows for correlations among the innovations of all included trend and cycle components. This flexible correlation structure has been demonstrated to be empirically important in many studies on unobserved components models (Morley et al., , 2003; Basistha and Nelson, , 2007; Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Hwu and Kim, , 2019).
A further contribution of this paper is the development of an MCMC estimation procedure and a marginal likelihood estimator to facilitate estimation and model comparison across various variants of the proposed SMUC-IV. Specifically, our proposed MCMC method builds upon the precision sampling algorithm for Gaussian state space models developed by Chan and Jeliazkov, (2009). A novel feature of our approach is that it exploits the joint Gaussianity between the observed data and the state parameters when deriving the conditional posterior of the latent states. This allows us to circumvent the need to apply Bayes’ rule, which is particularly cumbersome when the state and measurement equations are correlated (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Leiva-Leon and Uzeda, , 2023). Furthermore, one benefit of recognizing the joint Gaussianity between the observed data and the state parameters is that it permits direct derivation of the analytical expression for the likelihood function unconditional on the high-dimensional latent states, which is an essential component of our proposed marginal likelihood estimator. Specifically, in this paper we construct a conditional Monte Carlo improved modified harmonic mean estimator to compute the marginal likelihoods for various specifications of the SMUC-IV. The conditional Monte Carlo method was recently proposed by Chan, (2023) for estimating marginal likelihood for large vector autoregressions, and has been shown to significantly improve estimation accuracy.
In our empirical analysis, we jointly estimate four US macroeconomic stars—the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate. For the identification of the monetary policy shock, we employ the orthogonal high-frequency surprise series of Bauer and Swanson, (2023) as our external instrument for monetary policy shocks. The series expands the set of events to include Federal Reserve Chair speeches and orthogonalizes the surprises to pre-announcement variables to strengthen relevance and exogeneity.
To assess the validity of our framework, we conduct a Bayesian model comparison exercise. The results show that the proposed SMUC-IV is more strongly supported by the data than competing specifications. The findings support the relevance of the external instrument for monetary policy shocks, confirm correlations between long-run trends and short-run cycles, and underscore the significant influence of monetary policy shocks on the macroeconomic stars.
We document two main findings from our empirical analysis. First, contractionary monetary policy shocks lead to declines in the growth rate of potential output, trend inflation, and the neutral interest rate. The negative effects on the growth rate of potential output align with the main finding of Jordà et al., (2024) that monetary policy can have long-run effects on the real economy and implies an innovation channel, consistent with a growing literature showing that monetary policy can shape the economy’s long-run productive capacity through its effects on innovation and technological change (e.g., Stadler, , 1990; Moran and Queralto, , 2018; Ma and Zimmermann, , 2023; Fornaro and Wolf, , 2023; Meier and Reinelt, , 2024). Likewise, the decline in trend inflation is consistent with findings that contractionary monetary policy shocks lower long-term inflation expectations (e.g., Jarociński and Karadi, , 2020; Diegel and Nautz, , 2021), suggesting a re-anchoring channel when expectations drift from the inflation target. The decline in the neutral interest rate induced by contractionary monetary policy shocks may operate through two channels: the innovation channel and the re-anchoring channel. Contractionary shocks reduce the growth rate of potential output, and since it is a key driver of the neutral interest rate (see, e.g., Laubach and Williams, , 2003), this decline directly lowers the neutral interest rate. In addition, these shocks reduce trend inflation, which raises the real interest rate and encourages firms to reduce spending on R&D and productive investment—reinforcing the innovation channel and further depressing the neutral interest rate. Second, counterfactual analysis based on historical decompositions shows that, absent these contractionary shocks, the growth rate of potential output (along with the level of potential output), trend inflation, and the neutral interest rate would have been notably higher, implying that monetary policy shocks are important drivers of the stars. Taken together, the results imply that monetary policy can help re-anchor long-run inflation expectations at target, but over-tightening risks lowering the the growth rate of potential output and the neutral interest rate. Our two main findings are robust across a range of alternative specifications.
The paper is organized as follows. Section 2 introduces the proposed SMUC-IV. Section 3 details the prior distributions and develops an efficient MCMC method for estimation. Section 4 presents our modified harmonic mean estimator for marginal likelihood estimation, improved via Monte Carlo methods. Section 5 describes the data and empirical results, and Section 6 concludes.
2 Econometric Framework
In this section, we introduce our SMUC-IV. Section 2.1 presents the specification of the structural multivariate unobserved components model. Section 2.2 then outlines how the external instruments is employed for structural identification.
2.1 Model Specification
We consider the following trend and cycle decomposition:
$g_{t}$ $=g_{t}^{*}+c_{g,t},$ (1)
$\pi_{t}$ $=\pi_{t}^{*}+c_{\pi,t},$ (2)
$r_{t}$ $=\pi_{t}^{*}+r_{t}^{*}+c_{r,t},$ (3)
where $g_{t}$ denotes log level of real output, $\pi_{t}$ denotes inflation, and $r_{t}$ denotes the nominal interest rate. In addition, $g_{t}^{*}$, $\pi_{t}^{*}$, and $r_{t}^{*}$ represent the level of potential output, trend inflation, and the neutral interest rate, respectively. In this paper, we use the real GDP as the real output, GDP deflator inflation as the inflation, and the federal funds effective rate as the nominal interest rate (see Section 5.1 for more details). We assume inflation $\pi_{t}$ and nominal interest rate $r_{t}$ share a common trend component, $\pi_{t}^{*}$ as in Del Negro et al., (2017).
For the trend components, we assume the first difference of potential GDP follows a random walk process as in Grant and Chan, 2017b :
$\Delta g_{t}^{*}$ $=\Delta g_{t-1}^{*}+u_{t}^{g^{*}},$ (4)
where $\Delta g_{t}^{*}=g_{t}^{*}-g_{t-1}^{*}$ denotes the growth rate of potential output. For the trend inflation and neutral interest rate, we assume
$\pi_{t}^{*}$ $=\pi_{t-1}^{*}+u_{t}^{\pi^{*}},$ (5)
$r_{t}^{*}$ $=r_{t-1}^{*}+u_{t}^{r^{*}}.$ (6)
The initial conditions $g_{-1}^{*},g_{0}^{*},\pi_{0}^{*}$ and $r_{0}^{*}$ are treated as parameters to be estimated. Our trends are defined, consistent with the Beveridge-Nelson decomposition (Beveridge and Nelson, , 1981), as the infinite-horizon forecast of the actual variables of interest, conditional on the information set available in period $t$, which implies a random walk for the trends and stationary, mean-zero cycles.
The specifications of trend components broadly aligns with unobserved components models used to estimate macroeconomic stars, including strands that focus separately on the level of potential output (e.g., Grant and Chan, 2017a, ; Grant and Chan, 2017b, ), trend inflation (e.g., Chan et al., , 2013, 2018; Mertens, , 2016; Stock and Watson, , 2007, 2016; Hwu and Kim, , 2019; Eo et al., , 2023), and the neutral interest rate (e.g., Laubach and Williams, , 2003; Holston et al., , 2017; Del Negro et al., , 2017).
Regarding the cycle components, let $\mathbf{c}_{t}=(c_{g,t},c_{\pi,t},c_{r,t})^{\prime}$ be a vector of cycle components, and we assume that $\mathbf{c}_{t}$ evolves according to the following VAR($p$) process:
$\mathbf{c}_{t}=\bm{\Phi}_{1}\mathbf{c}_{t-1}+\cdots+\bm{\Phi}_{p}\mathbf{c}_{t-p}+\mathbf{u}_{t}^{c},$ (7)
where $\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p}$ are $3\times 3$ autoregressive coefficient matrices and $\mathbf{u}_{t}^{c}$ are the reduced-from innovations. [^1] For simplicity, we set the initial conditions $\mathbf{c}_{0}=\ldots=\mathbf{c}_{1-p}=\mathbf{0}$.
The state equations given in (4) - (6) can be expressed more compactly. To be specific, let $\bm{\tau}_{t}=(g_{t}^{*},\pi_{t}^{*},r_{t}^{*})^{\prime}$, we can rewrite (4) - (6) as
$\bm{\tau}_{t}=\bm{\Psi}_{1}\bm{\tau}_{t-1}+\bm{\Psi}_{2}\bm{\tau}_{t-2}+\mathbf{u}_{t}^{\tau},$ (8)
where $\bm{\Psi}_{1}=\text{diag}(2,1,1)$, $\bm{\Psi}_{2}=\text{diag}(-1,0,0)$ and $\mathbf{u}_{t}^{\tau}=(u_{t}^{g^{*}},u_{t}^{\pi^{*}},u_{t}^{r^{*}})^{\prime}$. Stacking equation (8) over equation (7), our model can be represented as
$\bm{\eta}_{t}=\mathbf{A}_{1}\bm{\eta}_{t-1}+\cdots+\mathbf{A}_{p}\bm{\eta}_{t-p}+\mathbf{u}_{t},$ (9)
where $\bm{\eta}_{t}=(\bm{\tau}_{t}^{\prime},\mathbf{c}_{t}^{\prime})^{\prime}$ is a vector containing the trend and cycle components. The coefficient matrices are defined as $\mathbf{A}_{1}=\text{diag}(\bm{\Psi}_{1},\bm{\Phi}_{1})$, $\mathbf{A}_{2}=\text{diag}(\bm{\Psi}_{2},\bm{\Phi}_{2})$, $\mathbf{A}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i})$ for $i=3,\ldots,p$. The residual vector $\mathbf{u}_{t}=(\mathbf{u}_{t}^{\tau^{\prime}},\mathbf{u}_{t}^{c^{\prime}})$ is of dimension $6\times 1$, which will be described shortly. To complete our model specification, we assume that the residual vector $\mathbf{u}_{t}$ are related to the structural shocks by
$\mathbf{u}_{t}=\mathbf{B}\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{6\times 1},\mathbf{I}_{6}),$ (10)
where $\bm{\epsilon}_{t}$ is a vector of structural shocks, $\mathbf{B}$ is the contemporaneous response matrix that is assumed to be non-singular, and $\bm{\Sigma}=\mathbf{B}\mathbf{B}^{\prime}$ is the covariance matrix of the residual vector $\mathbf{u}_{t}$. It is known that the contemporaneous response matrix $\mathbf{B}$, hence, the structural shocks $\bm{\epsilon}_{t}$, cannot be separately identified without additional information. In this paper, we employ external instrument to identify monetary policy shock. The next section provides the details of our identification approach.
Our proposed model specified in equations (9)–(10) provides a unified framework for studying the macroeconomic stars and their corresponding cycle components. Furthermore, we allows the innovations of all trend and cycle components, i.e., $\mathbf{u}_{t}$, to be correlated. This contrasts with conventional studies on unobserved components models (e.g., Watson, , 1986; Stock and Watson, , 2007; Chan et al., , 2018; Laubach and Williams, , 2003; Zaman, , 2025), which typically assume these innovations to be independent. Specifically, the contemporaneous response matrix $\mathbf{B}$ is assumed to be non-singular and unrestricted, which implies that $\bm{\Sigma}=\mathbf{B}\mathbf{B}^{\prime}$ is a full covariance matrix. In our empirical analysis, we assess the validity of this full correlation structure through a formal Bayesian model comparison exercise, which indicates strong evidence in favor of this modeling feature supported by the data.
2.2 Identification via External Instrument
In this section, we first discuss how the external instruments can be incorporated into our modeling framework for identification, and then describe our proposed SMUC-IV as an augmented strucutrual vector autoregression, a form which is used later for Bayesian estimation. In this paper, we consider the case in which one external instrument is used to identify one structural shock, namely, the monetary policy shock of interest. For the case of using multiple external instruments to identify multiple structural shocks, we refer readers to Arias et al., (2021), Braun and Brüggemann, (2023), and Hou, (2024) for more details.
To set the stage, we designate the last shock in $\bm{\epsilon}_{t}$, denoted by $\epsilon_{m,t}$, as the monetary policy shock of interest. Accordingly, we can write the structural shocks as $\bm{\epsilon}_{t}=(\bm{\epsilon}_{-m,t}^{\prime},\epsilon_{m,t}^{\prime})^{\prime}$ where $\bm{\epsilon}_{-m,t}$ is a vector containing all structural shocks other than $\epsilon_{m,t}$. Suppose that an external instrument $m_{t}$ is available, which is linked to the structural shocks $\bm{\epsilon}_{t}$ as
$m_{t}=\bm{\gamma}^{\prime}\bm{\epsilon}_{t}+\alpha v_{t},\quad v_{t}\sim\mathcal{N}(0,1),$ (11)
where $\bm{\gamma}$ is a $6\times 1$ vector of coefficient parameters associated with the structural shocks $\bm{\epsilon}_{t}$ and $v_{t}$ is the shock of the external instrument equation independent of $\bm{\epsilon}_{t}$, which can be interpreted as the measurement error of the external instrument.
To achieve identification of the monetary policy shock $\epsilon_{m,t}$, the external instrument $m_{t}$ is required to be correlated with $\epsilon_{m,t}$, but uncorrelated with $\bm{\epsilon}_{-m,t}$. More precisely, a valid external instrument needs to satisfy the following relevance and exogeneity conditions:
$\text{Relevance condition}:E(m_{t}\epsilon_{m,t})=\beta\neq 0,$
$\text{Exogeneity condition}:E(m_{t}\bm{\epsilon}_{-m,t}^{\prime})=\mathbf{0}_{1\times 5}.$
These two conditions are central to understanding how the external instrument can be used to identify the structural shock of interest $\epsilon_{m,t}$. Specifically, it conveys identifying information by distinguishing $\epsilon_{m,t}$ from the other shocks in $\bm{\epsilon}_{-m,t}$ through differences in their covariance structures with the the external instrument $m_{t}$.
The relevance and exogeneity conditions together provide further information by imposing zero restrictions on the parameter vector $\bm{\gamma}$. To see this, we first compute the covariance between $m_{t}$ and $\bm{\epsilon}_{t}$, that is
$E(m_{t}\bm{\epsilon}_{t}^{\prime})=\left(E(m_{t}\bm{\epsilon}_{-m,t}^{\prime}),E(m_{t}\epsilon_{m,t})\right)=(\mathbf{0}_{1\times 5},\beta)^{\prime},$
where the last equality is implied by the relevance and exogeneity conditions. On the other hand, given the external instrument equation (11), the covariance between $m_{t}$ and $\bm{\epsilon}_{t}$ can also be expressed as
$E(m_{t}\bm{\epsilon}_{t}^{\prime})=E\left((\bm{\gamma}^{\prime}\bm{\epsilon}_{t}+\alpha v_{t})\bm{\epsilon}_{t}^{\prime}\right)=\bm{\gamma}^{\prime},$
where the second equality holds because $v_{t}$ and $\bm{\epsilon}_{t}$ are assumed to be uncorrelated. Therefore, these results imply that
$\bm{\gamma}=(\mathbf{0}_{1\times 5},\beta)^{\prime}.$ (12)
To summarize, our SMUC-IV is specified as equations (9), (10), and (11), subject to the zero restrictions given in (12). The external instrument equation (11) with the zero restrictions in (12) gives $m_{t}=\beta\epsilon_{m,t}+\alpha v_{t}$, which is consistent with the specification used in Caldara and Herbst, (2019).
Our proposed SMUC-IV can be represented more compactly as an augmented structural vector autoregression for $\widetilde{\bm{\eta}}_{t}=(\bm{\eta}_{t}^{\prime},m_{t})^{\prime}$:
$\widetilde{\bm{\eta}}_{t}=\widetilde{\mathbf{A}}_{1}\widetilde{\bm{\eta}}_{t-1}+\cdots+\widetilde{\mathbf{A}}_{p}\widetilde{\bm{\eta}}_{t-p}+\widetilde{\mathbf{B}}\widetilde{\bm{\epsilon}}_{t},\quad\widetilde{\bm{\epsilon}}_{t}\sim\mathcal{N}(\mathbf{0}_{7\times 1},\mathbf{I}_{7}),$ (13)
where $\widetilde{\bm{\epsilon}}_{t}=(\bm{\epsilon}_{t}^{\prime},v_{t})^{\prime}$. The parameter matrices are given by
$\widetilde{\mathbf{A}}_{i}$ $=\begin{pmatrix}\mathbf{A}_{i}&\mathbf{0}_{6\times 1}\\ \mathbf{0}_{1\times 6}&0\end{pmatrix}\text{ for $i=1,\ldots,p$},$ (14)
$\widetilde{\mathbf{B}}$ $=\begin{pmatrix}\mathbf{B}&\mathbf{0}_{6\times 1}\\ \bm{\gamma}^{\prime}&\alpha\end{pmatrix}\text{ with $\bm{\gamma}=(\mathbf{0}_{1\times 5},\beta)^{\prime}$}.$ (15)
This augmented structural vector autoregression representation is commonly used in Bayesian analysis to facilitate posterior inference.
As indicated by recent studies (Arias et al., , 2021; Braun and Brüggemann, , 2023; Hou, , 2024), the use of external instruments can only achieve set-identification. For instance, in our case, it can be shown that the second to last column of $\widetilde{\mathbf{B}}$ that embeds the impact responses of monetary policy shocks, i.e., the last column of $\mathbf{B}$, which can only be identified up to sign changes. A common solution to this set-identification issue is to impose sign restrictions on the impulse responses, which are informed by economic theory. In this paper, rather than imposing dogmatic restrictions on response directions, we identify monetary policy shocks using an informative prior that accommodates estimation uncertainty. We will discuss this more in Section 3.1.
3 Bayesian Estimation
3.1 Priors
This section describes the prior distributions assigned to the model parameters. Let $\bar{\bm{\tau}}_{0}=(g_{-1}^{*},g_{0}^{*},\pi_{0}^{*},r_{0}^{*})^{\prime}$ denote the vector of the initial state parameters. Let $\bm{\Phi}_{l,i,j}$ represent the $(i,j)$ element of the autoregressive coefficient matrix $\bm{\Phi}_{l}$ for $i,j=1,2,3$, $l=1,\ldots,p$, and let $\mathbf{B}_{i,j}$ denote the $(i,j)$ element of the contemporaneous impact matrix $\mathbf{B}$ for $i,j=1,\ldots,6$. We assume the following independent priors:
$\bm{\Phi}_{l,i,j}\sim\mathcal{N}(\phi_{l,i,j},V_{\phi,l,i,j}),\quad\mathbf{B}_{i,j}\sim\mathcal{N}(b_{i,j},V_{b}),\quad\bar{\bm{\tau}}_{0}\sim\mathcal{N}(\bar{\bm{\tau}}_{00},\mathbf{V}_{\bar{\tau}_{00}}),$
$\beta\sim\mathcal{N}(\beta_{0},V_{\beta}),\quad\alpha\sim\mathcal{N}(\alpha_{0},V_{\alpha})\mathbf{1}(\alpha>0).$ (16)
Moreover, we consider a Minnesota-type adaptive hierarchical shrinkage prior for the autoregressive coefficients to address overfitting concerns in our richly parametrized model. To be specific, we set the prior mean of the autoregressive coefficient to $\phi_{l,i,j}=0$ and the prior variance to
$\mathbf{V}_{\phi,l,i,j}=\begin{dcases}\frac{\kappa_{1}}{l^{2}},\quad\text{$i=j$, $i,j=1,2,3$, $l=1,\ldots,p$},\\ \frac{\kappa_{2}\sigma^{2}_{i}}{l^{2}\sigma^{2}_{j}},\quad\text{$i\neq j$, $i,j=1,2,3$, $l=1,\ldots,p$}.\end{dcases}$
This specification reflects the prior belief that the coefficients on more distant lags are less important than those on recent lags, and are therefore shrunk more strongly toward zero. Following standard practice, the scale parameter $\sigma^{2}_{i}$ is set equal to the residual variance of an AR($p$) model for its corresponding variable $i$. The hyperparameters $\kappa_{1}$ and $\kappa_{2}$ control the shrinkage strength for the own-lag and cross-variable-lag coefficients, respectively. Empirical evidence from recent studies, such as Cross et al., (2020) and Chan, (2021), indicates that allowing the hyperparameters of shrinkage prior to be parameters to be estimated can substantially improves both forecasting accuracy and model fit. Therefore, rather than fixing these hyperparameters at predetermined values, we treat $\kappa_{1}$ and $\kappa_{2}$ as unknown parameters and assign them the following prior distributions:
$\kappa_{1}\sim\mathcal{U}(0,1),\quad\kappa_{2}\sim\mathcal{U}(0,1).$
We use an informative prior for the contemporaneous impact matrix $\mathbf{B}$, setting the prior variance to $V_{b}=0.01$ and specifying the prior mean as follows: [^2] In Section 5.6, we conduct a robustness analysis by treating $V_{b}$ as a unknown parameter to be estimated and the main finding our empirical results remain unchanged.
$b_{i,j}=\begin{dcases}0.1,\quad\text{$i=j$, $i,j=1,2,3$},\\ 1,\quad\text{$i=j$, $i,j=4,5,6$},\\ 0,\quad\text{$i\neq j$, $i,j=1,\ldots,6$.}\end{dcases}$
This prior centers the contemporaneous impact matrix $\mathbf{B}$ on a diagonal matrix, implying a priori uncorrelatedness among all reduced-form residuals for the cycle and trend components, with the prior standard deviations set to $0.1$ for the trend components and $1$ for the cycle components. [^3] Informative priors are commonly used in the estimation of multivariate unobserved component models. By imposing small standard deviations on the latent state parameters, this prior belief helps prevent overfitting and yields smoother, more economically sensible estimates.
The prior we consider here provides additional information for identifying the contemporaneous response matrix $\mathbf{B}$. Recall that the monetary policy shock $\epsilon_{m,t}$ is ordered last in $\bm{\epsilon}_{t}$. That means, the last column of $\mathbf{B}$ is the impact responses of the trend and cycle components to a monetary policy shock. Our prior reflects a belief that a one-standard-deviation increase in monetary policy shock $\epsilon_{m,t}$ results in a $100$ basis point increase in the cycle of interest rate $c_{r,t}$, while having no effect on $\Delta g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$, on impact. By the definition of $c_{r,t}$ in (3), the impact response of $c_{r,t}$ to a monetary policy shock is entirely attributable to the increase in $r_{t}$. This identifying information from our prior is similar to conventional studies on identifying monetary policy shock by normalize the magnitude of the interest rate response to a positive value. [^4] For instance, Bauer and Swanson, (2023) normalize the interest rate response to a monetary policy shock by $25$ basis points, while Miranda-Agrippino and Ricco, (2021) use a normalization of $100$ basis points. Unlike these approaches, we do not fix the size of the response of interest rate to a monetary policy shock, instead, we impose a relatively tight prior that centers the response at $100$ basis points. We also highlight that since we use an informative prior that centers the impact responses of $\Delta g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$ to a monetary policy shock at zeros, any nonzero effects found in our empirical study must be supported by the data.
For the parameters in the external instrument equation, we set $V_{\beta}=1$, $V_{\alpha}=1$, $\alpha_{0}=0$, and $\beta_{0}=0.5\times\sigma_{m}$, where $\sigma_{m}$ denotes the standard deviation of the external instrument. This is comparable to the setting in Caldara and Herbst, (2019). For the initial state parameters $\bar{\bm{\tau}}_{0}$, we use an uninformative prior by setting $\mathbf{V}_{\bar{\tau}_{00}}=100\mathbf{I}_{4}$ and $\bar{\bm{\tau}}_{00}=(g_{1},g_{1},\pi_{1},r_{1})^{\prime}$ where $g_{1}$, $\pi_{1}$ and $r_{1}$ are the first observations of log real GDP, inflation, and the interest rate in the sample period, respectively.
3.2 Posterior Sampler
We now discuss the estimation of our model with the prior described in the previous section. To set the stage, let $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, where $\mathbf{y}_{t}=(g_{t},\pi_{t},r_{t},m_{t})^{\prime}$, denote the vector of observed data, $\bm{\Phi}=(\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})$ denote the collection of the autoregressive coefficient matrices and $\bm{\tau}=(\bar{\bm{\tau}}_{0}^{\prime},\bm{\tau}_{1}^{\prime},\ldots,\bm{\tau}_{T}^{\prime})^{\prime}$ denote a vector of state parameters. The joint posterior distribution can be simulated by sequentially sampling from the following conditional distributions:
  1. 1.
    $p(\bm{\tau}|\mathbf{B},\beta,\alpha,\bm{\Phi},\kappa_{1},\kappa_{2},\mathbf{y})$;
  2. 2.
    $p(\mathbf{B},\beta,\alpha|\bm{\tau},\bm{\Phi},\kappa_{1},\kappa_{2},\mathbf{y})$;
  3. 3.
    $p(\bm{\Phi}|\bm{\tau},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2},\mathbf{y})$;
  4. 4.
    $p(\kappa_{1}|\bm{\tau},\mathbf{B},\beta,\alpha,\bm{\Phi},\kappa_{2},\mathbf{y})$;
  5. 5.
    $p(\kappa_{2}|\bm{\tau},\mathbf{B},\beta,\alpha,\bm{\Phi},\kappa_{1},\mathbf{y})$.
In Step 1, we sample the state vector $\bm{\tau}$ using the precision sampling method (Chan and Jeliazkov, , 2009; McCausland et al., , 2011; Rue, , 2001) instead of traditional Kalman filter techniques. The precision-based sampling method has gain increasing popularity for estimating state space models due to its computational efficiency and straightforward implementation. One complication under our framework, however, involves determining the conditional posterior distribution for $\bm{\tau}$. The conventional approach applies Bayes’ rule, which requires separately deriving both the conditional likelihood and the prior density of the state parameters. However, when the state and measurement equations are correlated, the derivation of the conditional posterior of the state parameters becomes more cumbersome (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Leiva-Leon and Uzeda, , 2023).
In this paper, we introduce a novel and direct approach to derive the conditional posterior of $\bm{\tau}$ that circumvents the use of Bayes’ rule. The key idea is to first obtain the joint conditional distribution of $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$, which will be shown later to be a Gaussian distribution, and then exploit the standard Gaussian conditioning properties to obtain the conditional posterior of $\bm{\tau}$. Specifically, by stacking equation (13) over $t=1,\ldots,T$, we obtain
[…]
where $\widetilde{\bm{\eta}}=(\widetilde{\bm{\eta}}_{1}^{\prime},\ldots,\widetilde{\bm{\eta}}_{T}^{\prime})^{\prime}$, $\widetilde{\bm{\Sigma}}=\widetilde{\mathbf{B}}\widetilde{\mathbf{B}}$,
$\mathbf{H}_{1}=\begin{pmatrix}\mathbf{I}_{7}&\mathbf{0}_{7}&\mathbf{0}_{7}&\cdots&\cdots&\cdots&\cdots&\mathbf{0}_{7}\\ -\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{p}&\ddots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\vdots\\ \mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\mathbf{0}_{7}\\ \mathbf{0}_{7}&\cdots&\mathbf{0}_{7}&-\widetilde{\mathbf{A}}_{p}&\cdots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}\end{pmatrix},\quad\bm{\Xi}=\begin{pmatrix}\bm{\Xi}_{1}\\ \mathbf{0}_{4\times 4}\\ \bm{\Xi}_{2}\\ \mathbf{0}_{4\times 4}\\ \mathbf{0}_{7\times 4}\\ \vdots\\ \mathbf{0}_{7\times 4}\end{pmatrix}$
with
$\bm{\Xi}_{1}=\begin{pmatrix}-1&2&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix},\quad\bm{\Xi}_{2}=\begin{pmatrix}0&-1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}$.
[…]
$\mathbf{H}_{2}\begin{pmatrix}\bar{\bm{\tau}}_{0}\\ \widetilde{\bm{\eta}}\end{pmatrix}=\widetilde{\bm{\tau}}+\mathbf{e},\quad\mathbf{e}\sim\mathcal{N}\left(\mathbf{0},\bm{\Omega}\right),$ (18)
[…]
$\mathbf{H}_{2}=\begin{pmatrix}\mathbf{I}_{4}&\mathbf{0}_{4\times 7T}\\ -\bm{\Xi}&\mathbf{H}_{1}\end{pmatrix}$.
Based on the result in (18), we now derive the joint conditional distribution of $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$. First, by definition, it can be verified that $\widetilde{\bm{\eta}}_{t}$ relates linearly to $(\bm{\tau}_{t}^{\prime},\mathbf{y}_{t}^{\prime})^{\prime}$ as:
$\widetilde{\bm{\eta}}_{t}=\mathbf{Q}\begin{pmatrix}\bm{\tau}_{t}\\ \mathbf{y}_{t}\end{pmatrix}$, (19)
for $t=1,\ldots,T$, where
$\mathbf{Q}=\begin{pmatrix}\mathbf{I}_{3}&\mathbf{0}_{3}&\mathbf{0}_{3\times 1}\\ \widetilde{\mathbf{Q}}&\mathbf{I}_{3}&\mathbf{0}_{3\times 1}\\ \mathbf{0}_{1\times 3}&\mathbf{0}_{1\times 3}&1\end{pmatrix},\quad\widetilde{\mathbf{Q}}=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&-1&-1\end{pmatrix}$.
Substituting the linear relationship in (19) into (18) yields
$\mathbf{H}\mathbf{z}=\widetilde{\bm{\tau}}+\mathbf{e},\quad\mathbf{e}\sim\mathcal{N}\left(\mathbf{0},\bm{\Omega}\right),$
where $\mathbf{z}=(\bar{\bm{\tau}}_{0}^{\prime},\bm{\tau}_{1}^{\prime},\mathbf{y}_{1}^{\prime},\ldots,\bm{\tau}_{T}^{\prime},\mathbf{y}_{T}^{\prime})^{\prime}$ and
$\mathbf{H}=\mathbf{H}_{2}\begin{pmatrix}\mathbf{I}_{4}&\mathbf{0}_{4\times 7T}\\ \mathbf{0}_{7T\times 4}&\mathbf{I}_{T}\otimes\mathbf{Q}\end{pmatrix}$.
This suggests that $\mathbf{z}\sim\mathcal{N}(\bm{\mu}_{z},\mathbf{K}_{z}^{-1})$ with $\bm{\mu}_{z}=\mathbf{H}^{-1}\widetilde{\bm{\tau}}$ and $\mathbf{K}_{z}=\mathbf{H}^{\prime}\bm{\Omega}^{-1}\mathbf{H}$.
Note that since $\mathbf{z}$ has a Gaussian distribution and contains the same elements as $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$ up to permutation, the standard properties of Gaussian distribution imply that $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$ is also Gaussian. More precisely, let $\mathbf{P}_{z}$ be a permutation matrix such that $\mathbf{P}_{z}\mathbf{z}=(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$. Then we have the following result:
$\left((\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}|\mathbf{B},\beta,\alpha,\bm{\Phi},\kappa_{1},\kappa_{2}\right)\sim\mathcal{N}(\bm{\mu},\mathbf{K}^{-1}),$ (20)
where $\bm{\mu}=\mathbf{P}_{z}\bm{\mu}_{z}$ and $\mathbf{K}=\mathbf{P}_{z}\mathbf{K}_{z}\mathbf{P}_{z}^{\prime}$. Finally, we partition the mean vector $\bm{\mu}$ and precision matrix $\mathbf{K}$ according to the dimensions of $\bm{\tau}$ and $\mathbf{y}$ as
$\bm{\mu}=\begin{pmatrix}\bm{\mu}_{\tau}\\ \bm{\mu}_{y}\end{pmatrix},\quad\mathbf{K}=\begin{pmatrix}\mathbf{K}_{\tau}&\mathbf{K}_{\tau,y}\\ \mathbf{K}_{\tau,y}^{\prime}&\mathbf{K}_{y}\end{pmatrix}$.
By applying the standard Gaussian conditioning results, the conditional posterior of $\bm{\tau}$ is given by
$(\bm{\tau}|\mathbf{B},\beta,\alpha,\bm{\Phi},\kappa_{1},\kappa_{2},\mathbf{y})\sim\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1}),$
where $\widehat{\bm{\tau}}=\bm{\mu}_{\tau}-\mathbf{K}_{\tau}^{-1}\mathbf{K}_{\tau,y}(\mathbf{y}-\bm{\mu}_{y})$. It can be check that the precision matrix $\mathbf{K}_{\tau}$ is a banded matrix. Therefore, one can efficiently sample from $\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1})$ using the precision-based approach of Chan and Jeliazkov, (2009).
For sampling $(\mathbf{B},\beta,\alpha)$ in Step 2, we first note that these parameters correspond to the nonzero elements in the contemporaneous impact matrix $\widetilde{\mathbf{B}}$ of the augmented structural autoregression representation in (13). Consequently, sampling $(\mathbf{B},\beta,\alpha)$ is equivalent to sampling $\widetilde{\mathbf{B}}$ with the zero restrictions in (15) and the sign restriction on $\alpha$. Since these restrictions can be expressed as linear equality and inequality restrictions on $\widetilde{\mathbf{B}}$, we sample the nonzero parameters using the efficient sampler developed by Hou, (2024). We refer readers to this paper for more implementation details.
Since Step 3 - Step 5 of the posterior sampler implement either standard Bayesian procedures or minor modifications of established methods, we relegate their technical details to Appendix A.
4 Marginal Likelihood Estimation
Marginal likelihood is the standard criterion for Bayesian model comparison. In this section, we present an approach for estimating the marginal likelihood of our proposed SMUC-IV and its restricted versions. Our method builds upon the modified harmonic mean estimator proposed by Gelfand and Dey, (1994), integrating it with the conditional Monte Carlo method of Chan, (2023) to enhance estimation accuracy. We begin with an overview of the modified harmonic mean estimator and then outline our conditional Monte Carlo improved estimator for the SMUC-IV. Technical details are relegated to Appendix B.
4.1 Modified Harmonic Mean Estimator
The marginal likelihood of a given model is defined as
$p(\mathbf{y})=\int p(\mathbf{y}|\bm{\theta})p(\bm{\theta})d\bm{\theta},$
where $\mathbf{y}$ denotes a vector of observed data, $\bm{\theta}$ denotes the set of all parameters specific to the model, $p(\mathbf{y}|\bm{\theta})$ is the likelihood function, and $p(\bm{\theta})$ is the prior density for the model. The modified harmonic mean estimator is built upon the following identity:
$p(\mathbf{y})^{-1}=\int\frac{q(\bm{\theta})}{p(\mathbf{y}|\bm{\theta})p(\bm{\theta})}p(\bm{\theta}|\mathbf{y})d\bm{\theta}.$
Here $p(\bm{\theta}|\mathbf{y})$ is the posterior distribution and $q(\bm{\theta})$ is a tuning function that can be any density function defined on $\bm{\theta}$ with its support contained in the support of the posterior density, i.e., $q(\bm{\theta})>0$ implies $p(\bm{\theta}|\mathbf{y})>0$. This suggests that the marginal likelihood can be estimated using the following estimator:
$\widehat{p(\mathbf{y})}_{GD}=\left(\frac{1}{R}\sum_{i=1}^{R}\frac{q(\bm{\theta}^{(i)})}{p(\mathbf{y}|\bm{\theta}^{(i)})p(\bm{\theta}^{(i)})}\right)^{-1},$
where $\bm{\theta}^{(1)},\ldots,\bm{\theta}^{(R)}$ are draws from the posterior distribution $p(\bm{\theta}|\mathbf{y})$.
While the modified harmonic mean estimator described above is simulation-consistent and straightforward to implement, it might perform poorly when the model parameter $\bm{\theta}$ is high-dimensional. To improve estimation accuracy, we employ the conditional Monte Carlo method of Chan, (2023). The key idea is to first analytically integrate out as many parameters in $\bm{\theta}$ as possible, and then construct a modified harmonic mean estimator using the resulting unconditional likelihood and prior on the remaining lower-dimensional parameters.
4.2 Estimating the Marginal Likelihood for SMUC-IV
We now outline our conditional Monte Carlo improved modified harmonic mean estimator for the marginal likelihood of the SMUC-IV. In our modeling setting, the marginal likelihood is given by
$p(\mathbf{y})$ $=\int p(\mathbf{y}|\bm{\tau},\bm{\Phi},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2})p(\bm{\tau},\bm{\Phi},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2})d(\bm{\tau},\bm{\Phi},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2})$
$=\int p(\mathbf{y}|\bm{\tau},\bm{\Phi},\mathbf{B},\alpha,\beta)p(\bm{\tau}|\bm{\Phi},\mathbf{B},\alpha,\beta)p(\bm{\Phi}|\kappa_{1},\kappa_{2})p(\kappa_{1})p(\kappa_{2})p(\mathbf{B})p(\alpha)p(\beta)d(\bm{\tau},\bm{\Phi},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2})$
$=\int p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)p(\mathbf{B})p(\bm{\Phi})p(\alpha)p(\beta)d(\bm{\Phi},\mathbf{B},\beta,\alpha).$
The second equality follows from the conditional independence structure of the prior distributions. In the last equality, we have integrated out the state parameters $\bm{\tau}$ and hyperparameters $(\kappa_{1},\kappa_{2})$. Specifically, we have:
$p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)=\int p(\mathbf{y}|\bm{\tau},\bm{\Phi},\mathbf{B},\alpha,\beta)p(\bm{\tau}|\bm{\Phi},\mathbf{B},\alpha,\beta)d\bm{\tau},$ (21)
$p(\bm{\Phi})=\int p(\bm{\Phi}|\kappa_{1},\kappa_{2})p(\kappa_{1})p(\kappa_{2})d(\kappa_{1},\kappa_{2}).$ (22)
Note that the expression in (21) can be obtained directly from result given in (20), which shows that the joint conditional distribution of $(\mathbf{y},\bm{\tau})$ is Gaussian. Consequently, $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$ can be derived using standard properties of the Gaussian distribution. See Appendix B for more details.
Using the analytical expressions for (21) and (22), we can estimate the marginal likelihood of our SMUC-IV with the following conditional Monte Carlo improved modified harmonic mean estimator:
$\widehat{p(\mathbf{y})}_{CMGD}=\left(\frac{1}{R}\sum_{i=1}^{R}\frac{q(\bm{\Phi}^{(i)},\mathbf{B}^{(i)},\alpha^{(i)},\beta^{(i)})}{p(\mathbf{y}|\bm{\Phi}^{(i)},\mathbf{B}^{(i)},\alpha^{(i)},\beta^{(i)})p(\mathbf{B}^{(i)})p(\bm{\Phi}^{(i)})p(\alpha^{(i)})p(\beta^{(i)})}\right)^{-1},$ (23)
where $(\bm{\Phi}^{(i)},\mathbf{B}^{(i)},\alpha^{(i)},\beta^{(i)})$, $i=1,\ldots,R$, are posterior draws that can be obtained using the posterior sampler described in the last section. Note that the key difference between this estimator and the standard modified harmonic mean estimator is the use of the likelihood function $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$, which is unconditional on the high-dimensional state parameters $\bm{\tau}$, and the marginal prior $p(\bm{\Phi})$. This formulation substantially reduces the dimensionality of the numerical integration, which in turn reduces the Monte Carlo variance of the estimator, resulting in greater numerical stability and estimation precision (see Chan, (2023) for more discussion).
We now turn to the choice of the tuning density function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$, a critical component of our estimator (23) that is essential for the accuracy of marginal likelihood estimation. Although the posterior density is the theoretically optimal choice, it is computationally intractable. Therefore, we follow Chan, (2023) and approximate the posterior using a truncated Gaussian density. Specifically, we consider a tuning density function that takes the following form:
$q(\bm{\Phi},\mathbf{B},\alpha,\beta)=q_{\Phi}(\bm{\Phi})q_{B}(\mathbf{B})q_{\alpha}(\alpha)q_{\beta}(\beta),$ (24)
where each $q_{j}(\cdot)$ for $j\in\{\Phi,B,\alpha,\beta\}$ is a truncated Gaussian density with mean and covariance matrix set to the estimated posterior mean and covariance matrix. [^5] Using an appropriately truncated tuning density can ensure that the modified harmonic mean estimator has a finite variance (Geweke, , 1999). The full details for $q_{\Phi}(\bm{\Phi})$, $q_{B}(\mathbf{B})$, $q_{\alpha}(\alpha)$, and $q_{\beta}(\beta)$ are provided in Appendix B.
5 Empirical Analysis
In this section, we first describe the data in Section 5.1 and then conduct a model comparison exercise in Section 5.2 to validate our proposed SMUC-IV framework. Section 5.3 presents our estimates of the macroeconomic stars—namely, the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate. Section 5.4 investigates the effects of monetary policy shocks on these stars. Subsequently, Section 5.5 assesses the role of the monetary policy shocks in driving the evolution of the macroeconomic stars. Finally, robustness analyses are reported in Section 5.6.
5.1 Data and External Instrument
We estimate the empirical model using the following quarterly data from 1987:Q4 to 2023:Q4: real GDP, the GDP deflator, the federal funds effective rate, the shadow interest rate, and the orthogonal monetary policy surprise (Bauer and Swanson, , 2023). To capture both conventional and unconventional monetary policy when the nominal federal funds rate is at the effective lower bound (ELB), we use the Wu and Xia, (2016) shadow short rate. Consistent with the quarterly frequency of our data we use four lags in our model. We transform real GDP as $100\times\log(x_{t})$ and compute the quarterly GDP deflator as annualized log growth, $400\times\log(x_{t}/x_{t-1})$. The federal funds effective rate, shadow interest rate, and orthogonal monetary policy surprise are originally monthly series; we obtain their quarterly versions by averaging the three monthly observations within each quarter. Real GDP, the GDP deflator, and the federal funds effective rate are available from the FRED database. We consider different measures of inflation and the interest rate as well as consider the bond premium (Favara et al., , 2016) as a control variable in the vector autoregression, see section 5.6. The shadow interest rate is from the website of the Federal Reserve Bank of Atlanta, the excess bond premium is from the website of the Board of Governors of the Federal Reserve System, and the orthogonal monetary policy surprise series is from the website of the Federal Reserve Bank of San Francisco.
In this paper, we use the orthogonal monetary policy surprise series developed by Bauer and Swanson, (2023) as our external instrument. This series identifies monetary policy shocks from high-frequency asset price movements in narrow windows around policy announcements. High-frequency interest rate changes around FOMC announcements are a common tool for identifying monetary policy effects, but recent studies have questioned their exogeneity and relevance as instruments, particularly for estimating macroeconomic impacts (e.g., Ramey, , 2016; Miranda-Agrippino and Ricco, , 2021). For instance, monetary policy surprises may be correlated with publicly available macroeconomic and financial data released before FOMC announcements. To address these concerns, Bauer and Swanson, (2023) (i) expand the set of monetary policy events to include speeches by the Federal Reserve Chair—roughly doubling the number of announcements—and (ii) orthogonalize the resulting surprises with respect to pre-announcement macroeconomic and financial data to account for predictability via the “Fed response to news” channel.
5.2 Model Comparison Exercise
This section conducts a Bayesian model comparison exercise by evaluating the marginal likelihoods of alternative specifications of our SMUC-IV.
The first alternative specification imposes the restriction that monetary policy shocks have no contemporaneous effects on all of the trend components, that is, the macroeconomic stars of interest in this paper. We refer to this model as SMUC-IV-R1. Next, we consider a restricted version of our SMUC-IV that imposes zero correlation between the external instrument and all structural shocks by setting $\beta=0$. A comparison between our proposed SMUC-IV and SMUC-IV-R2 serves to test the relevance condition of the instrument. Note that under the restriction $\beta=0$, the proxy equation (11) is independent of the system of equations in the unobserved components model specified in (9). Therefore, by imposing various patterns of zero restrictions on $\mathbf{B}$, we can assess different correlation structures between the trend and cycle innovations. In our model comparison exercise, we further consider two nested versions of SMUC-IV-R2. The first nested version of SMUC-IV-R2 is denoted as SMUC-IV-R3. This specification assumes that innovations between the trend and cycle components are uncorrelated, while allowing for correlation within the innovations of the trend components and within the innovations of the cycle components, respectively. Specifically, under SMUC-IV-R3 we assume the $6\times 6$ contemporaneous response matrix $\mathbf{B}$ to be a block diagonal matrix with each block of dimension $3\times 3$. For the second nested version of SMUC-IV-R2, we assume $\mathbf{B}$ to be a diagonal matrix and denote this model as SMUC-IV-R4. This specification is similar to that in many conventional studies on multivariate unobserved components models, which assumes all innovations of the trend and cycle components are mutually independent. Table 1 provides a summary of the competing models in our model comparison exercise.
Table 1: Competing models used in the comparison exercise.
Model Description
SMUC-IV The baseline model specified in equations (13) - (15).
SMUC-IV-R1 A restricted version of SMUC-IV by imposing zero contemporaneous effects of monetary policy shocks on the trend components.
SMUC-IV-R2 A restricted version of the SMUC-IV that imposes zero correlation between the external instrument and all structural shocks by setting $\beta=0$.
SMUC-IV-R3 A restricted version of the SMUC-IV-R2 that imposes zero correlation between trend and cycle innovations
SMUC-IV-R4 A restricted version of the SMUC-IV-R2 that imposes zero correlation between all trend and cycle innovations.
Table 2 reports the log marginal likelihood estimates. The results provide significant evidence that our proposed SMUC-IV model outperforms all alternative specifications considered. A few findings are also worth highlighting. First, by comparing the SMUC-IV with the SMUC-IV-R1, the difference in the log marginal likelihoods is about $13$, strongly supporting the non-zero contemporaneous effects of monetary policy shocks on the trend components. Second, the SMUC-IV-R2 performs better than the SMUC-IV-R3 and SMUC-IV-R4. This suggests that allowing for correlated innovations among all trend and cycle components is an important modeling feature supported by the data. In particular, the log marginal likelihood of SMUC-IV-R2 is about $24$ higher than that of SMUC-IV-R3, indicating that the assumption of independent innovations for the trend and cycle components is empirically implausible in our empirical analysis. Lastly, the SMUC-IV outperforms SMUC-IV-R1, suggesting that the relevance condition of the instrument is satisfied, although the difference in the log marginal likelihoods between these two models is only about 33.666In Section 5.6, we also conduct a robustness analysis by considering an alternative instrument in our application, and we find that our main empirical results remain robust under this alternative choice of instrument.
Table 2: Estimated log marginal likelihoods
SMUC-IV SMUC-IV-R1 SMUC-IV-R2 SMUC-IV-R3 SMUC-IV-R4
-207 -220 -210 -223 -231
5.3 Estimates of Macroeconomic Stars
Although macroeconomic stars—for example, the neutral interest rate—have been defined in various ways across the literature, most theoretical frameworks suggest that shifts in the low-frequency components of macroeconomic variables are closely linked to movements in any theoretically defined star. Accordingly, we examine the persistent evolution of macroeconomic stars through the lens of trend dynamics. Using US data, we jointly estimate four stars: the level of potential GDP, $g^{*}$, the growth rate of potential GDP, $\Delta g^{*}$, and trend inflation, $\pi^{*}$, and the neutral interest rate, $r^{*}$. These stars are measured as long-run trends in the spirit of the Beveridge-Nelson decomposition (Beveridge and Nelson, , 1981).
[…]
Panel (a) of Figure 1 presents the estimates of $^{*}. gg^{*}$ rises over time but experiences two notable drops during the 2007–2009 Global Financial Crisis and the COVID-19 pandemic. In both cases, the level of $g^{*}$ fails to return to its pre-crisis trajectory, suggesting that transitory shocks may translate into permanent reductions in the long-run growth path.
Panel (b) of Figure 1 shows estimates of $\Delta g^{*}$. From the early to late 1990s, $\Delta g^{*}$ rises noticeably, likely reflecting the internet technology boom. After the 1990s, however, it follows a downward trend until around 2010. Then it follows an upward trend until the early 2020s. Our estimates show a similar tendency to those in (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Maffei-Faccioli, , 2025), albeit with considerable fluctuations. Sharp declines are evident during major crises such as the dot-com bust in the early 2000s, the Global Financial Crisis of 2007–2009, and the COVID-19 pandemic. These episodes coincide with well-documented hysteresis effects, where severe recessions leave lasting scars on the economy’s productive capacity (Cerra and Saxena, , 2005, 2008; Cerra et al., , 2023).
Panel (c) of Figure 1 reports the estimates of $\pi^{*}$. Consistent with prior work (e.g., Stock and Watson, , 2007, 2016; Chan et al., , 2018; Eo et al., , 2023), $\pi^{*}$ displays remarkable stability: after falling in the 1990s, it remains anchored near 2% from the 2000s until the pandemic period. This pattern highlights the Federal Reserve’s success in stabilizing long-run inflation expectations.
Finally, Panel (d) of Figure 1 presents estimates of $r^{*}$. The estimates show a steady decline from the 1990s onward, followed by relative stability during the 2010s-2020s. Our estimates of $r^{*}$ capture the secular decline in the neutral interest rate documented in the literature on estimating $r^{*}$ (e.g., Laubach and Williams, , 2003; Lubik and Matthes, , 2015; Del Negro et al., , 2017; Morley et al., , 2024).
Figure 1: The plot compares the actual data with trend estimates along 68% credible bands.
5.4 Do Monetary Policy Shocks Affect Macroeconomic Stars?
Having obtained plausible estimates of the macroeconomic stars, we now turn to the paper’s central aim: exploring the effects of monetary policy shocks on these stars. Specifically, we examine how monetary policy influences $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$. Monetary policy shocks are identified using an external instrument based on high-frequency policy surprises developed by Bauer and Swanson, (2023).
Figure 2: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. Note that the responses are constant over time.
Figure 2 shows the posterior distributions of the impulse responses of $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$ to a one-standard-deviation contractionary monetary policy shock. This shock raises the median response of the interest rate by about 0.29 basis points on impact (see Figure 3). By construction, the responses are constant over time, as the model assumes the stars follow a random walk.
Panel (a) of Figure 2 shows that $\Delta g^{*}$ falls by about 0.1 percentage point. The negative effects on the growth rate of potential output align with the main finding of Jordà et al., (2024) that monetary policy can have long-run effects on the real economy, suggesting an innovation channel whereby monetary tightening curbs spendings on R&D and innovation, limits productivity gains, and reduces productive investment, thereby depressing the growth rate of potential output (see, e.g., Stadler, , 1990; Moran and Queralto, , 2018; Ma and Zimmermann, , 2023; Fornaro and Wolf, , 2023; Meier and Reinelt, , 2024).
Panel (b) of Figure 2 shows that contractionary monetary policy shocks also reduce $\pi^{*}$ by about 0.2 percentage points. At first glance, significant effects of monetary policy shocks on trend inflation may appear undesirable, as they seem to imply de-anchoring of long-run inflation expectations. However, when long-run expectations are persistently off target, such effects may reflect a re-anchoring channel, in which policy deliberately shifts expectations back toward the target. As shown by our counterfactual analysis (see Figure 4), monetary policy shocks pull trend inflation down make it move around the 2% target from mid-1990s to the onset of Covid-19 pandemic. This finding is consistent with evidence that such shocks lower survey-based long-run inflation expectations, pointing to a re-anchoring channel in which monetary policy shifts long-run inflation expectations back toward target when they drift persistently away from it (see, e.g., Jarociński and Karadi, , 2020; Diegel and Nautz, , 2021).
Panel (c) shows that $r^{*}$ falls by about 0.3 basis points. This decline, caused by contractionary monetary policy shocks, may operate through two complementary mechanisms. First, the innovation channel in panel (a) indicates that contractionary monetary policy shocks lowers potential output growth; because trend growth is a key determinant of the neutral rate (e.g., Laubach and Williams, , 2003), this directly pushes $r^{*}$ down. Second, the re-anchoring channel in panel (b) indicates that these shocks lower $\pi^{*}$, thereby raising the real interest rate and also encouraging firms to cut spending on R&D and productive investment —reinforcing the innovation channel and further depressing $r^{*}$.
A more practical question concerns how actual macroeconomic variables respond to contractionary monetary policy shocks once policy-induced shifts in the macroeconomic stars are taken into account. Figure 3 shows that the impulse responses of the actual variables align with standard macroeconomic predictions: the nominal interest rate rises, output falls, and inflation declines. Panel (a) of Figure 3 shows that output falls by about 1.5 percent after five years. Panel (b) indicates that inflation declines by roughly 0.17 percentage points after five years, following the 0.29-basis-point increase in the nominal interest rate shown on impact in Panel (c). Notably, when monetary policy shocks are allowed to affect the macroeconomic stars, the responses of the actual variables become markedly more persistent.
Figure 3: The plots shows the impulse responses with 68% impulse responses of the actual variable to a monetary policy shock.
5.5 Are the Monetary Policy Shocks a Key Driver for Shaping the Macroeconomic Stars?
Given the meaningful effects of monetary policy shocks on the macroeconomic stars, an important question is whether these shocks have played a important role in shaping the historical dynamics of the stars. To address this, we conduct a counterfactual analysis based on the historical decomposition of the trend components, as shown in Figure 4.
Figure 4: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
Figure 4 compares the actual trend paths with counterfactual paths, which are generated by shutting down all identified monetary policy shocks, following the methodology outlined in Kilian and Lütkepohl, (2017). The shaded regions represent 68% credible bands around the counterfactual estimates.
Our counterfactual analysis quantifies the contributions of monetary policy shocks over time. For instance, in Figure 4, Panel (a) shows that, absent these shocks, $\Delta g^{*}$ would have been over 1 percent higher during the crisis; Panel (b) indicates that $g^{*}$ would have been persistently higher, with a cumulative gap of about 0.5 percent over the sample period. Panel (c) shows that $\pi^{*}$, which remained below its counterfactual path since the early 1990s, would have hovered closer to 3 percent in the absence of repeated monetary tightening. Panel (d) shows that $r^{*}$ would have been more than 2 basis points higher during the financial crisis.
Taken together, the counterfactual analysis emphasizes that monetary policy is a main driver of the macroeconomic stars and tends to play a dual role: it effectively stabilizes long-run inflation expectations but can also impose a persistent drag on the real economy’s long-run trends when over-tightening occurs. Policymakers should therefore weigh the benefits of anchoring inflation against the potential costs for the real economy in the long run.
5.6 Robustness Analysis
We assess the robustness of our baseline findings by re-estimating the model under a range of alternative specifications. Figures C.1-C.8 show results of the IRFs of the trends and C.9-C.16 show the historical decomposition of all alternative specifications. Across all variations, the qualitative responses of the macroeconomic stars to a contractionary monetary policy shock remain similar to those in the baseline, indicating that our main results are not driven by specific modelling choices.
First, estimating the shrinkage parameter $V_{b}$ via a hierarchical Bayes approach yields posterior impulse responses that closely match the baseline, with slightly wider credible intervals (Figure C.1) and very similar historical decompositions (Figure C.9). Second, restricting the sample to pre-COVID observations ending in 2019Q4 produces similar negative responses of $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$ (Figure C.2) as well as an similar counterfactual path of the trends (Figure C.10), suggesting that the pandemic period does not drive the results. Third, increasing the VAR lag length from four to eight (Figure C.3 and igure C.11) does not materially alter the estimated responses or historical decompositions. Fourth, replacing the federal funds rate with the one-year or two-year Treasury yield (Figures C.4, C.5, C.12 and C.13) preserves the qualitative pattern of the responses and the pattern of the counterfactual pathes. Fifth, substituting the GDP deflator with the PCE price index (Figures C.6 and C.14) yields comparable estimates. Sixth, using the unadjusted monetary policy surprise measure from Bauer and Swanson (2023) instead of the orthogonalized version (Figures C.7 and C.15) results in nearly identical posterior distributions. Finally, controlling for financial conditions by adding the excess bond premium (Favara et al., , 2016) to the cycle part of the model confirms our main empirical findings( Figures C.8 and C.16).
Overall, these robustness analyses confirm that the estimated effects of monetary policy shocks on the macroeconomic stars are stable across alternative sample periods, lag specifications, interest rate measures, inflation measures, external instruments, and controlling for financial conditions using the excess bond premium.
6 Conclusions
This paper develops SMUC-IV to explore the effects of monetary policy shocks on the macroeconomic stars, measured as long-run trends, in a unified framework. Using our SMUC-IV, we show that monetary tightening can have negative effects on the growth rate of potential output, trend inflation, and the neutral interest rate, and that monetary policy shocks are an important driver of the growth rate of potential output (along with the level of potential output), trend inflation, and the neutral interest rate. From a policy perspective, the results indicate that policymakers can use monetary policy to re-anchor long-run inflation expectations at target, but also caution that too much tightening could harm the real economy in the long run. Recently, growing attention has been paid to the long-run effects not only of monetary policy shocks but also of other structural shocks, such as financial, fiscal, and demand shocks (Cerra and Saxena, , 2005, 2008; Antolin-Diaz and Surico, , 2025; Furlanetto et al., , 2025). Against this backdrop, future research could extend our framework to examine how different types of structural shocks shape long-run macroeconomic equilibria.
Appendix Appendix A Estimation Details
This Appendix provides details of Step 3 - Step 5 of the posterior sampler.
To sample $\bm{\Phi}$ in Step 3, we first rewrite the model in (13) as
[…]
where $\bar{\bm{\eta}}_{t}=(\Delta g_{t}^{*}-\Delta g_{t-1}^{*},\pi_{t}^{*}-\pi_{t-1}^{*},r_{t}^{*}-r_{t-1}^{*},\mathbf{c}_{t}^{\prime},m_{t})^{\prime}$, $\widetilde{\bm{\Sigma}}=\widetilde{\mathbf{B}}\widetilde{\mathbf{B}}^{\prime}$ and $\bar{\mathbf{A}}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i},0)$ for $i=1,\ldots,p$. Stacking equation (25) over $t=1,\ldots,T$, we have
[…]
where $\bar{\bm{\eta}}=(\bar{\bm{\eta}}_{1}^{\prime},\ldots,\bar{\bm{\eta}}_{T}^{\prime})^{\prime}$, $\bar{\mathbf{a}}=\text{vec}\left((\bar{\mathbf{A}}_{1},\ldots,\bar{\mathbf{A}}_{p})^{\prime}\right)$, $\bar{\mathbf{X}}=(\bar{\mathbf{X}}_{1}^{\prime},\ldots,\bar{\mathbf{X}}_{T}^{\prime})^{\prime}$ with $\bar{\mathbf{X}}_{t}=\mathbf{I}_{7}\otimes(\bar{\bm{\eta}}_{t-1}^{\prime},\ldots,\bar{\bm{\eta}}_{t-p}^{\prime})$. Let $\bm{\phi}=\text{vec}\left((\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})^{\prime}\right)$ and $\mathbf{S}_{\phi}$ be the selection matrix such that $\bar{\mathbf{a}}=\mathbf{S}_{\phi}\bm{\phi}$. Then we can write (26) as
[…]
where $\mathbf{X}=\bar{\mathbf{X}}\mathbf{S}_{\phi}$. Since we consider a Gaussian prior $\bm{\phi}\sim\mathcal{N}(\bm{\phi}_{0},\mathbf{V}_{\phi_{0}})$, where the prior mean $\bm{\phi}_{0}$ and the diagonal covariance matrix $\mathbf{V}_{\phi_{0}}$ can be constructed as described in Section 3.1. Using the standard Bayesian linear regression results, we can obtain the conditional posterior
$(\bm{\phi}|\bm{\tau},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2},\mathbf{y})\sim\mathcal{N}(\widehat{\bm{\phi}},\widehat{\mathbf{V}}_{\phi}),$
where $\widehat{\mathbf{V}}_{\phi}=(\mathbf{X}^{\prime}(\mathbf{I}_{T}\otimes\widetilde{\bm{\Sigma}})\mathbf{X}+\mathbf{V}_{\phi_{0}}^{-1})^{-1}$ and $\widehat{\bm{\phi}}=\widehat{\mathbf{V}}_{\phi}(\mathbf{X}^{\prime}(\mathbf{I}_{T}\otimes\widetilde{\bm{\Sigma}})\bar{\bm{\eta}}+\mathbf{V}_{\phi_{0}}^{-1}\bm{\phi}_{0})$.
To obtain the conditional posteriors of $\kappa_{1}$ and $\kappa_{2}$ in Step 4 and Step 5, we first define
$\widetilde{\bm{\Phi}}_{l,i,j}$ $=\begin{dcases}l^{2}\left(\bm{\Phi}_{l,i,j}-\phi_{l,i,j}\right)^{2},\qquad\text{$l=1,\ldots,p$, $i,j=1,2,3$, $i=j$},\\ \frac{\sigma_{j}^{2}}{\sigma_{i}^{2}}l^{2}\left(\bm{\Phi}_{l,i,j}-\phi_{l,i,j}\right)^{2},\qquad\text{$l=1,\ldots,p$, $i,j=1,2,3$, $i\neq j$},\end{dcases}$
Then, it can be shown that the density functions of the conditional posteriors of $\kappa_{1}$ and $\kappa_{2}$ are given by
$p(\kappa_{1}|\bm{\tau},\mathbf{B},\beta,\alpha,\bm{\Phi},\mathbf{y})\propto\kappa_{1}^{-\frac{3p}{2}}e^{-\frac{1}{2\kappa_{1}}\sum_{i=j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}}\mathbf{1}(0<\kappa_{1}<1),$
$p(\kappa_{2}|\bm{\tau},\mathbf{B},\beta,\alpha,\bm{\Phi},\mathbf{y})\propto\kappa_{2}^{-\frac{6p}{2}}e^{-\frac{1}{2\kappa_{2}}\sum_{i\neq j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}}\mathbf{1}(0<\kappa_{2}<1).$
This implies that the conditional posteriors of $\kappa_{1}$ and $\kappa_{2}$ are truncated inverse-Gamma distributions:
$(\kappa_{1}|\bm{\tau},\mathbf{B},\beta,\alpha,\bm{\Phi},\mathbf{y})$ $\sim\mathcal{IG}\left(\frac{3p}{2}-1,\frac{1}{2}\sum_{i=j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}\right)\mathbf{1}(0<\kappa_{1}<1),$
$(\kappa_{2}|\bm{\tau},\mathbf{B},\beta,\alpha,\bm{\Phi},\mathbf{y})$ $\sim\mathcal{IG}\left(\frac{6p}{2}-1,\frac{1}{2}\sum_{i\neq j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}\right)\mathbf{1}(0<\kappa_{2}<1).$
Appendix Appendix B Details of Marginal Likelihood Estimation
In this Appendix, we first present the analytical expressions for (21) and (22) and then detail the construction of the tuning function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$.
Expressions for $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$ and $p(\bm{\Phi})$
From result in (20), let $\bm{\Lambda}=\mathbf{K}^{-1}$ and partition $\bm{\Lambda}$ according to the dimensions of $\bm{\tau}$ and $\mathbf{y}$ as
$\bm{\Lambda}=\begin{pmatrix}\bm{\Lambda}_{\tau}&\bm{\Lambda}_{\tau,y}\\ \bm{\Lambda}_{\tau,y}^{\prime}&\bm{\Lambda}_{y}\end{pmatrix}$.
The property of Gaussian distribution implies that the marginal distribution of $\mathbf{y}$ (of dimension $4T\times 1$) is a Gaussian with means $\bm{\mu}_{y}$ and covariance matrix $\bm{\Lambda}$. This yields
$p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)=(2\pi)^{-2T}|\bm{\Lambda}_{y}|^{-\frac{1}{2}}e^{-\frac{1}{2}\left(\mathbf{y}-\bm{\mu}_{y}\right)^{\prime}\bm{\Lambda}^{-1}_{y}\left(\mathbf{y}-\bm{\mu}_{y}\right)}.$
Next, we derive the expression for the marginal prior $p(\bm{\Phi})$. Given the priors of $\bm{\Phi}$ and $(\kappa_{1},\kappa_{2})$ described in Section 3.1, we have
$p(\bm{\Phi})$ $=\int p(\bm{\Phi}|\kappa_{1},\kappa_{2})p(\kappa_{1})p(\kappa_{2})d(\kappa_{1},\kappa_{2})$
$=c_{\kappa}\int\kappa_{1}^{-\frac{3p}{2}}\kappa_{2}^{-\frac{6p}{2}}e^{-\frac{1}{2\kappa_{1}}\sum_{i=j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}}e^{-\frac{1}{2\kappa_{2}}\sum_{i\neq j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}}\mathbf{1}(0<\kappa_{1}<1,0<\kappa_{2}<1)d(\kappa_{1},\kappa_{2})$
$=c_{\kappa}\left(\int_{0}^{1}\kappa_{1}^{-\frac{3p}{2}}e^{-\frac{1}{2\kappa_{1}}\sum_{i=j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}}d\kappa_{1}\right)\times\left(\int_{0}^{1}\kappa_{2}^{-\frac{6p}{2}}e^{-\frac{1}{2\kappa_{2}}\sum_{i\neq j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}}d\kappa_{2}\right)$
$=c_{\kappa}\times\Gamma(\hat{\nu}_{1})\widehat{S}_{1}^{-\hat{\nu}_{1}}F_{\mathcal{IG}}(1;\hat{\nu}_{1}\hat{S}_{1})\times\Gamma(\hat{\nu}_{2})\widehat{S}_{2}^{-\hat{\nu}_{2}}F_{\mathcal{IG}}(1;\hat{\nu}_{2}\hat{S}_{2}),$
where the normalizing constant $c_{\kappa}=(2\pi)^{-\frac{9p}{2}}\prod_{i=1}^{3}\prod_{j=1}^{3}\prod_{l=1}^{p}\frac{\sigma_{j}l}{\sigma_{i}}$, $\widetilde{\bm{\Phi}}_{l,i,j}$ is defined in Appendix A, $F_{\mathcal{IG}}(x;\nu,S)$ denotes the cumulative distribution function of an inverse-gamma distribution with shape parameter $\nu$ and scale parameter $S$ evaluated at $x$, and
$\hat{\nu}_{1}=\frac{3p}{2}-1,\qquad\hat{S}_{1}=\frac{1}{2}\sum_{i=j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j},\qquad\hat{\nu}_{2}=\frac{6p}{2}-1,\qquad\hat{S}_{2}=\frac{1}{2}\sum_{i\neq j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}.$
The last equality holds because the integrals in the third line are the kernels of inverse-gamma densities.
Details about the Tuning Density Function
This section provides the expressions for the truncated Gaussian densities $q_{\Phi}(\bm{\Phi})$, $q_{B}(\mathbf{B})$, $q_{\alpha}(\alpha)$, and $q_{\beta}(\beta)$, which are given as follows:
$q_{\Phi}(\bm{\Phi})$ $=c^{-1}_{\Phi}(2\pi)^{-\frac{9p}{2}}|\widehat{\bm{\Sigma}}_{\Phi}|^{-\frac{1}{2}}e^{-\frac{1}{2}\left(\text{vec}(\bm{\Phi})-\text{vec}(\widehat{\bm{\Phi}})\right)^{\prime}\widehat{\bm{\Sigma}}_{\Phi}^{-1}\left(\text{vec}(\bm{\Phi})-\text{vec}(\widehat{\bm{\Phi}})\right)}\mathbf{1}(\text{vec}(\bm{\Phi})\in\mathcal{R}_{\Phi}),$
$q_{B}(\mathbf{B})$ $=c^{-1}_{B}(2\pi)^{-18}|\widehat{\bm{\Sigma}}_{B}|^{-\frac{1}{2}}e^{-\frac{1}{2}\left(\text{vec}(\mathbf{B})-\text{vec}(\widehat{\mathbf{B}})\right)^{\prime}\widehat{\bm{\Sigma}}_{B}^{-1}\left(\text{vec}(\mathbf{B})-\text{vec}(\widehat{\mathbf{B}})\right)}\mathbf{1}(\text{vec}(\mathbf{B})\in\mathcal{R}_{B}),$
$q_{\beta}(\beta)$ $=c_{\beta}^{-1}(2\pi\widehat{\sigma}^{2}_{\beta})^{-\frac{1}{2}}e^{-\frac{1}{2\widehat{\sigma}^{2}_{\beta}}(\beta-\widehat{\beta})^{2}}\mathbf{1}(\beta\in\mathcal{R}_{\beta}),$
$q_{\alpha}(\alpha)$ $=c_{\alpha}^{-1}(2\pi\widehat{\sigma}^{2}_{\alpha})^{-\frac{1}{2}}e^{-\frac{1}{2\widehat{\sigma}^{2}_{\alpha}}(\alpha-\widehat{\alpha})^{2}}\mathbf{1}(\alpha\in\mathcal{R}_{\alpha}),$
where $c_{\Phi}$, $c_{B}$, $c_{\beta}$, and $c_{\alpha}$ are normalizing constants. The parameters $\widehat{\bm{\Phi}}$, $\widehat{\mathbf{B}}$, $\widehat{\beta}$, and $\widehat{\alpha}$ are the estimated posterior means, while $\widehat{\bm{\Sigma}}_{\Phi}$, $\widehat{\bm{\Sigma}}_{B}$, $\widehat{\sigma}^{2}_{\beta}$, and $\widehat{\sigma}^{2}_{\alpha}$ are their corresponding estimated covariance matrices and variances. The truncation regions are given by:
$\mathcal{R}_{\Phi}$ $=\left\{\bm{\Phi}:\left(\text{vec}(\bm{\Phi})-\text{vec}(\widehat{\bm{\Phi}})\right)^{\prime}\widehat{\bm{\Sigma}}_{\Phi}^{-1}\left(\text{vec}(\bm{\Phi})-\text{vec}(\widehat{\bm{\Phi}})\right)<\chi^{2}_{0.95,9p}\right\},$
$\mathcal{R}_{B}$ $=\left\{\mathbf{B}:\left(\text{vec}(\mathbf{B})-\text{vec}(\widehat{\mathbf{B}})\right)^{\prime}\widehat{\bm{\Sigma}}_{B}^{-1}\left(\text{vec}(\mathbf{B})-\text{vec}(\widehat{\mathbf{B}})\right)<\chi^{2}_{0.95,36}\right\},$
$\mathcal{R}_{\beta}$ $=\left\{\beta:\frac{(\beta-\widehat{\beta})^{2}}{\widehat{\sigma}^{2}_{\beta}}<\chi^{2}_{0.95,1}\right\},$
$\mathcal{R}_{\alpha}$ $=(0,w),$
where $\chi^{2}_{k_{1},k_{2}}$ denotes the $k_{1}$ quantile of the chi-square distribution with $k_{2}$ degrees of freedom. The upper bound $w$ for $\mathcal{R}_{\alpha}$ is chosen such that the probability mass of the untruncated Gaussian $\mathcal{N}(\widehat{\alpha},\widehat{\sigma}^{2}_{\alpha})$ over the interval $(0,w)$ is 0.95. Formally, $w$ satisfies:
$\mathcal{N}_{cdf}\left(\frac{w-\widehat{\alpha}}{\widehat{\sigma}_{\alpha}}\right)-\mathcal{N}_{cdf}\left(-\frac{\widehat{\alpha}}{\widehat{\sigma}_{\alpha}}\right)=0.95,$
where $\mathcal{N}_{cdf}(\cdot)$ is the cumulative distribution function of the standard normal distribution.
Appendix Appendix C Results Robustness Checks
Figure C.1: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the shrinkage parameter $V_{b}$ using a hierachical Bayes approach.
Figure C.2: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we use only data unitl 2019Q4.
Figure C.3: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with eight lags instead of four.
Figure C.4: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with the one-year treasury yield instead of using the federal funds effective rate.
Figure C.5: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with the two-year treasury yield instead of using the federal funds effective rate.
Figure C.6: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with the personal consumption expenditure price index instead of the GDP deflator.
Figure C.7: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (Bauer and Swanson, , 2023) instead of using the orthogonalized monetary policy surprise measure.
Figure C.8: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with adding the excess bond premium.
Figure C.9: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the shrinkage parameter $\lambda_{B}$ using a hierachical Bayes approach.
Figure C.10: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we use only data unitl 2019Q4.
Figure C.11: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with eight lags instead of four.
Figure C.12: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with the one-year treasury yield instead of using the federal funds effective rate.
Figure C.13: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with the two-year treasury yield instead of using the federal funds effective rate.
Figure C.14: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with the personal consumption expenditure price index instead of the GDP deflator.
Figure C.15: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (Bauer and Swanson, , 2023) instead of using the orthogonalized monetary policy surprise measure.
Figure C.16: The plot compares the actual trends and the counterfactual path with 68% credible bands. The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017). In this plot we estimate the model with adding the excess bond premium.

### v2

@@
-Assessing the Effects of Monetary Shocks on Macroeconomic Stars: A SMUC-IV Framework
+
 Abstract
 This paper proposes a structural multivariate unobserved components model with external instrument (SMUC-IV) to investigate the effects of monetary policy shocks on key U.S.
-macroeconomic “stars”—namely, the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate.
-A key feature of our approach is the use of an external instrument to identify monetary policy shocks within the multivariate unobserved components modeling framework.
+macroeconomic “stars”—namely, the potential GDP growth, trend inflation, and the neutral interest rate.
+A key feature of our approach is the use of an external instrument to identify monetary policy shocks within the multivariate unobserved components modelling framework.
 We develop an MCMC estimation method to facilitate posterior inference within our proposed SMUC-IV framework.
-In addition, we propose an marginal likelihood estimator to enable model comparison across alternative specifications.
-Our empirical analysis shows that contractionary monetary policy shocks have significant negative effects on the macroeconomic stars, highlighting the nonzero long-run effects of transitory monetary policy shocks.
+In addition, we propose a marginal likelihood estimator to enable model comparison across alternative specifications.
+Our empirical analysis shows that contractionary monetary policy shocks have negative effects on the macroeconomic stars, highlighting the non-zero long-run effects of transitory monetary policy shocks.
 Keywords: macroeconomic stars, monetary policy, unobserved components models, external instrument, precision sampling, marginal likelihood estimation
@@
 1 Introduction
-Long-run equilibrium levels of key macroeconomic variables—commonly referred to as “stars”—are central to guiding countercyclical policy and assessing long-run performance.
-In the conventional view, the macroeconomic stars are unaffected by monetary policy shocks, which are thought to have only transitory effects.
-Recently, a growing body of evidence challenges this view, showing that monetary shocks can have long-run influences on the economy (e.g., Hanson and Stein, , 2015; Diegel and Nautz, , 2021; Jordà et al., , 2024).
+Long-run equilibrium levels of key macroeconomic variables—commonly referred to as “stars”—are central for guiding countercyclical policy and assessing long-run performance.
+In the conventional view, the macroeconomic stars are unaffected by monetary policy shocks.
+These shocks are typically believed to have only transitory effects.
+However, recent evidence challenges this view.
+For instance, jorda2024long show that contractionary monetary policy shocks have negative and long-lasting effects on output.
 A less studied empirical question, however, is how monetary policy shocks affect the macroeconomic stars.
+This question is of interest because it provides guidance on how policymakers can bring inflation back to target while supporting sustainable economic growth in the wake of the heightened inflation following the COVID-19 pandemic.
 The main objective of this paper is to develop a unified econometric framework to address this question.
-Multivariate unobserved components models have become the standard econometric tool for analyzing macroeconomic stars (e.g., Kuttner, , 1994; Laubach and Williams, , 2003; Chan et al., , 2016; Zaman, , 2025).
+Multivariate unobserved components models have become the standard econometric tool for analyzing macroeconomic stars (e.g., kuttner1994estimating; laubach2003measuring; chan2016bounded; zaman2025unified).
 These models decompose observed macroeconomic series into long-run trends and short-run cycles, identifying the stars with the trend components.
@@
 The main novelty of our approach is the use of an external instrument to identify monetary policy shocks for structural analysis within a multivariate unobserved components model, which we refer to as SMUC-IV.
-While the use of external instruments to identify macroeconomic shocks in structural vector autoregressions has become increasingly common (Mertens and Ravn, , 2013; Ramey, , 2016; Stock and Watson, , 2018; Jarociński and Karadi, , 2020; Caldara and Herbst, , 2019; Arias et al., , 2021), very few studies have considered their application within multivariate unobserved components models.
+While the use of external instruments to identify macroeconomic shocks in structural vector autoregressions has become increasingly common (mertens2013dynamic; ramey2016macroeconomic; stock2018identification; jarocinski2020deconstructing; caldara2019monetary; arias2021inference), very few studies have considered their application within multivariate unobserved components models.
 The SMUC-IV introduced in this paper addresses this gap.
@@
 Our proposed SMUC-IV possesses two desirable characteristics.
-First, it offers a unified framework for the joint estimation of the macroeconomic stars and the assessment of their responses to monetary policy shocks.
+First, it offers a unified framework that jointly estimates the macroeconomic stars and assesses their responses to monetary policy shocks, rather than focusing on a single star.
 This ensures internally consistent estimates of the macroeconomic stars, enabling a coherent analysis of the dynamic interactions between the long-run trends and short-run cyclical fluctuations.
 Second, our SMUC-IV specification allows for correlations among the innovations of all included trend and cycle components.
-This flexible correlation structure has been demonstrated to be empirically important in many studies on unobserved components models (Morley et al., , 2003; Basistha and Nelson, , 2007; Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Hwu and Kim, , 2019).
-A further contribution of this paper is the development of an MCMC estimation procedure and a marginal likelihood estimator to facilitate estimation and model comparison across various variants of the proposed SMUC-IV.
-Specifically, our proposed MCMC method builds upon the precision sampling algorithm for Gaussian state space models developed by Chan and Jeliazkov, (2009).
-A novel feature of our approach is that it exploits the joint Gaussianity between the observed data and the state parameters when deriving the conditional posterior of the latent states.
-This allows us to circumvent the need to apply Bayes’ rule, which is particularly cumbersome when the state and measurement equations are correlated (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Leiva-Leon and Uzeda, , 2023).
-Furthermore, one benefit of recognizing the joint Gaussianity between the observed data and the state parameters is that it permits direct derivation of the analytical expression for the likelihood function unconditional on the high-dimensional latent states, which is an essential component of our proposed marginal likelihood estimator.
-Specifically, in this paper we construct a conditional Monte Carlo improved modified harmonic mean estimator to compute the marginal likelihoods for various specifications of the SMUC-IV.
-The conditional Monte Carlo method was recently proposed by Chan, (2023) for estimating marginal likelihood for large vector autoregressions, and has been shown to significantly improve estimation accuracy.
-In our empirical analysis, we jointly estimate four US macroeconomic stars—the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate.
-For the identification of the monetary policy shock, we employ the orthogonal high-frequency surprise series of Bauer and Swanson, (2023) as our external instrument for monetary policy shocks.
+This flexible correlation structure has been shown to be empirically important in many studies on unobserved components models (morley2003beveridge; basistha2007new; grant2017bayesian; grant2017reconciling; hwu2019estimating).
+A further contribution of this paper is the development of an MCMC estimation procedure and a marginal likelihood estimator, designed to facilitate estimation and model comparison across various variants of the proposed SMUC-IV.
+Specifically, our proposed MCMC method builds upon the precision sampling algorithm for Gaussian state space models developed by chan2009efficient.
+A novel feature of our approach is that it exploits the joint Gaussianity between the observed data and the state parameters, which allows for the direct derivation of the conditional posterior of the states.
+This approach bypasses the need for separate computations of the conditional likelihood and prior, leading to a more straightforward implementation, particularly when the state and measurement equations are correlated (grant2017bayesian; grant2017reconciling; leiva2023endogenous).
+A related benefit of recognizing the joint Gaussianity between the observed data and the state parameters is that it permits direct derivation of the likelihood function unconditional on the high-dimensional latent states.
+This unconditional likelihood is an essential component for our proposed marginal likelihood estimator.
+Specifically, we construct a conditional Monte Carlo improved modified harmonic mean estimator to compute the marginal likelihoods for various specifications of the SMUC-IV.
+The conditional Monte Carlo method was recently proposed by chan2023comparing for estimating marginal likelihood for large vector autoregressions, and has been shown to significantly improve estimation accuracy.
+In our empirical analysis, we jointly estimate three US macroeconomic stars—the potential GDP growth, trend inflation, and the neutral interest rate.
+For the identification of the monetary policy shock, we employ the orthogonal high-frequency surprise series of bauer2023reassessment as our external instrument for monetary policy shocks.
 The series expands the set of events to include Federal Reserve Chair speeches and orthogonalizes the surprises to pre-announcement variables to strengthen relevance and exogeneity.
-To assess the validity of our framework, we conduct a Bayesian model comparison exercise.
+To assess the validity of our framework, we conduct a Bayesian model comparison exercise in our empirical analysis.
 The results show that the proposed SMUC-IV is more strongly supported by the data than competing specifications.
-The findings support the relevance of the external instrument for monetary policy shocks, confirm correlations between long-run trends and short-run cycles, and underscore the significant influence of monetary policy shocks on the macroeconomic stars.
-We document two main findings from our empirical analysis.
-First, contractionary monetary policy shocks lead to declines in the growth rate of potential output, trend inflation, and the neutral interest rate.
-The negative effects on the growth rate of potential output align with the main finding of Jordà et al., (2024) that monetary policy can have long-run effects on the real economy and implies an innovation channel, consistent with a growing literature showing that monetary policy can shape the economy’s long-run productive capacity through its effects on innovation and technological change (e.g., Stadler, , 1990; Moran and Queralto, , 2018; Ma and Zimmermann, , 2023; Fornaro and Wolf, , 2023; Meier and Reinelt, , 2024).
-Likewise, the decline in trend inflation is consistent with findings that contractionary monetary policy shocks lower long-term inflation expectations (e.g., Jarociński and Karadi, , 2020; Diegel and Nautz, , 2021), suggesting a re-anchoring channel when expectations drift from the inflation target.
-The decline in the neutral interest rate induced by contractionary monetary policy shocks may operate through two channels: the innovation channel and the re-anchoring channel.
-Contractionary shocks reduce the growth rate of potential output, and since it is a key driver of the neutral interest rate (see, e.g., Laubach and Williams, , 2003), this decline directly lowers the neutral interest rate.
-In addition, these shocks reduce trend inflation, which raises the real interest rate and encourages firms to reduce spending on R&D and productive investment—reinforcing the innovation channel and further depressing the neutral interest rate.
-Second, counterfactual analysis based on historical decompositions shows that, absent these contractionary shocks, the growth rate of potential output (along with the level of potential output), trend inflation, and the neutral interest rate would have been notably higher, implying that monetary policy shocks are important drivers of the stars.
-Taken together, the results imply that monetary policy can help re-anchor long-run inflation expectations at target, but over-tightening risks lowering the the growth rate of potential output and the neutral interest rate.
+The findings support the relevance of the external instrument for monetary policy shocks, confirm correlations between long-run trends and short-run cycles, and—most importantly—underscore the influence of monetary policy shocks on the macroeconomic stars.
+We document two main findings on how monetary policy shocks affect the macroeconomic stars.
+First, contractionary monetary policy shocks lead to declines in the macroeconomic stars.
+The negative effects on potential GDP growth are consistent with the empirical evidence in moran2018innovation, garga2021output, ma2023monetary, jorda2024long, and meier2024monetary, which shows that tighter monetary policy dampens investment in innovation and, in turn, productivity.
+Likewise, the decline in the neutral interest rate following contractionary monetary policy shocks may be interpreted as reflecting the same adverse effects of monetary tightening on investment in innovation.
+The fall in trend inflation is consistent with findings that contractionary monetary policy shocks lower long-term inflation expectations (e.g., jarocinski2020deconstructing; diegel2021long).
+Second, counterfactual analysis based on historical decompositions shows that, in the absence of these contractionary shocks, potential GDP growth, trend inflation, and the neutral interest rate would have been notably higher, implying that monetary policy is a driving force in shaping the stars.
 Our two main findings are robust across a range of alternative specifications.
+The negative effects of contractionary monetary policy shocks on potential GDP growth and the neutral interest rate can be rationalized by the mechanism highlighted in the Keynesian endogenous growth model with nominal rigidities in fornaro2023scars, which shows that monetary tightening may trigger a “supply–demand doom loop”: lower demand reduces firms’ profits, which curtails investment; weaker investment slows expected productivity growth, lowers household wealth, and further depresses demand—ultimately leading to substantial declines in potential GDP growth and the neutral interest rate.
+The fall in trend inflation may reflect a re-anchoring channel, whereby expectations that have drifted away from the inflation target are deliberately pulled back, as discussed in detail in diegel2021long.
+The results imply that central banks face a dilemma: disinflating the economy may come at the cost of lower potential GDP growth and a lower neutral interest rate.
+This, in turn, points to a potential role for fiscal interventions that support business investment and the economy’s productive capacity during periods of disinflation.
 The paper is organized as follows.
@@
 Section 2.1 presents the specification of the structural multivariate unobserved components model.
-Section 2.2 then outlines how the external instruments is employed for structural identification.
+Section 2.2 then outlines how the external instrument is employed for structural identification.
 2.1 Model Specification
@@
 $\pi_{t}$ $=\pi_{t}^{*}+c_{\pi,t},$ (2)
-$r_{t}$ $=\pi_{t}^{*}+r_{t}^{*}+c_{r,t},$ (3)
-where $g_{t}$ denotes log level of real output, $\pi_{t}$ denotes inflation, and $r_{t}$ denotes the nominal interest rate.
-In addition, $g_{t}^{*}$, $\pi_{t}^{*}$, and $r_{t}^{*}$ represent the level of potential output, trend inflation, and the neutral interest rate, respectively.
-In this paper, we use the real GDP as the real output, GDP deflator inflation as the inflation, and the federal funds effective rate as the nominal interest rate (see Section 5.1 for more details).
-We assume inflation $\pi_{t}$ and nominal interest rate $r_{t}$ share a common trend component, $\pi_{t}^{*}$ as in Del Negro et al., (2017).
-For the trend components, we assume the first difference of potential GDP follows a random walk process as in Grant and Chan, 2017b :
-$\Delta g_{t}^{*}$ $=\Delta g_{t-1}^{*}+u_{t}^{g^{*}},$ (4)
-where $\Delta g_{t}^{*}=g_{t}^{*}-g_{t-1}^{*}$ denotes the growth rate of potential output. For the trend inflation and neutral interest rate, we assume
+$i_{t}$ $=\pi_{t}^{*}+r_{t}^{*}+c_{i,t},$ (3)
+where $g_{t}$ denotes real GDP growth, $\pi_{t}$ denotes inflation, and $i_{t}$ denotes the nominal interest rate.
+In addition, $g_{t}^{*}$, $\pi_{t}^{*}$, and $r_{t}^{*}$ represent the potential GDP growth, trend inflation, and the neutral interest rate, respectively.
+In this paper, we use the log difference of real GDP as the real GDP growth, GDP deflator inflation as the inflation, and the federal funds effective rate as the nominal interest rate (see Section 5.1 for more details).
+We assume inflation $\pi_{t}$ and nominal interest rate $i_{t}$ share a common trend component, $\pi_{t}^{*}$ as in del2017safety.
+We assume the trend components of real GDP growth, inflation and real interest rate to follow random walk processes:
+$g_{t}^{*}$ $=g_{t-1}^{*}+u_{t}^{g^{*}},$ (4)
 $\pi_{t}^{*}$ $=\pi_{t-1}^{*}+u_{t}^{\pi^{*}},$ (5)
 $r_{t}^{*}$ $=r_{t-1}^{*}+u_{t}^{r^{*}}.$ (6)
-The initial conditions $g_{-1}^{*},g_{0}^{*},\pi_{0}^{*}$ and $r_{0}^{*}$ are treated as parameters to be estimated.
-Our trends are defined, consistent with the Beveridge-Nelson decomposition (Beveridge and Nelson, , 1981), as the infinite-horizon forecast of the actual variables of interest, conditional on the information set available in period $t$, which implies a random walk for the trends and stationary, mean-zero cycles.
-The specifications of trend components broadly aligns with unobserved components models used to estimate macroeconomic stars, including strands that focus separately on the level of potential output (e.g., Grant and Chan, 2017a, ; Grant and Chan, 2017b, ), trend inflation (e.g., Chan et al., , 2013, 2018; Mertens, , 2016; Stock and Watson, , 2007, 2016; Hwu and Kim, , 2019; Eo et al., , 2023), and the neutral interest rate (e.g., Laubach and Williams, , 2003; Holston et al., , 2017; Del Negro et al., , 2017).
-Regarding the cycle components, let $\mathbf{c}_{t}=(c_{g,t},c_{\pi,t},c_{r,t})^{\prime}$ be a vector of cycle components, and we assume that $\mathbf{c}_{t}$ evolves according to the following VAR($p$) process:
+The initial conditions $g_{0}^{*},\pi_{0}^{*}$ and $r_{0}^{*}$ are treated as parameters to be estimated.
+Our trends are defined, consistent with the Beveridge-Nelson decomposition (beveridge1981new), as the infinite-horizon forecast of the actual variables of interest, conditional on the information set available in period $t$, which implies a random walk for the trends and stationary, mean-zero cycles.
+The specifications of trend components broadly aligns with unobserved components models used to estimate macroeconomic stars, including strands that focus separately on the potential GDP growth (e.g., grant2017bayesian; grant2017reconciling), trend inflation (e.g., chan2013new; chan2018new; mertens2016measuring; stockwatson2; stock2016core; hwu2019estimating; eo2023understanding), and the neutral interest rate (e.g., laubach2003measuring; holston2017measuring; del2017safety).
+Regarding the cycle components, let $\mathbf{c}_{t}=(c_{g,t},c_{\pi,t},c_{i,t})^{\prime}$ be a vector of cycle components, and we assume that $\mathbf{c}_{t}$ evolves according to the following VAR($p$) process:
 $\mathbf{c}_{t}=\bm{\Phi}_{1}\mathbf{c}_{t-1}+\cdots+\bm{\Phi}_{p}\mathbf{c}_{t-p}+\mathbf{u}_{t}^{c},$ (7)
-where $\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p}$ are $3\times 3$ autoregressive coefficient matrices and $\mathbf{u}_{t}^{c}$ are the reduced-from innovations.
+where $\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p}$ are $3\times 3$ autoregressive coefficient matrices and $\mathbf{u}_{t}^{c}$ are the reduced-form innovations.
 [^1] For simplicity, we set the initial conditions $\mathbf{c}_{0}=\ldots=\mathbf{c}_{1-p}=\mathbf{0}$.
 The state equations given in (4) - (6) can be expressed more compactly. To be specific, let $\bm{\tau}_{t}=(g_{t}^{*},\pi_{t}^{*},r_{t}^{*})^{\prime}$, we can rewrite (4) - (6) as
-$\bm{\tau}_{t}=\bm{\Psi}_{1}\bm{\tau}_{t-1}+\bm{\Psi}_{2}\bm{\tau}_{t-2}+\mathbf{u}_{t}^{\tau},$ (8)
-where $\bm{\Psi}_{1}=\text{diag}(2,1,1)$, $\bm{\Psi}_{2}=\text{diag}(-1,0,0)$ and $\mathbf{u}_{t}^{\tau}=(u_{t}^{g^{*}},u_{t}^{\pi^{*}},u_{t}^{r^{*}})^{\prime}$.
-Stacking equation (8) over equation (7), our model can be represented as
+$\bm{\tau}_{t}=\bm{\tau}_{t-1}+\mathbf{u}_{t}^{\tau},$ (8)
+where $\mathbf{u}_{t}^{\tau}=(u_{t}^{g^{*}},u_{t}^{\pi^{*}},u_{t}^{r^{*}})^{\prime}$. Stacking equation (8) over equation (7), our model can be represented as
 $\bm{\eta}_{t}=\mathbf{A}_{1}\bm{\eta}_{t-1}+\cdots+\mathbf{A}_{p}\bm{\eta}_{t-p}+\mathbf{u}_{t},$ (9)
 where $\bm{\eta}_{t}=(\bm{\tau}_{t}^{\prime},\mathbf{c}_{t}^{\prime})^{\prime}$ is a vector containing the trend and cycle components.
-The coefficient matrices are defined as $\mathbf{A}_{1}=\text{diag}(\bm{\Psi}_{1},\bm{\Phi}_{1})$, $\mathbf{A}_{2}=\text{diag}(\bm{\Psi}_{2},\bm{\Phi}_{2})$, $\mathbf{A}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i})$ for $i=3,\ldots,p$.
+The coefficient matrices are $\mathbf{A}_{1}=\text{diag}(\mathbf{I}_{3},\bm{\Phi}_{1})$, $\mathbf{A}_{i}=\text{diag}(\mathbf{0}_{3\times 3},\bm{\Phi}_{i})$ for $i=2,\ldots,p$.
 The residual vector $\mathbf{u}_{t}=(\mathbf{u}_{t}^{\tau^{\prime}},\mathbf{u}_{t}^{c^{\prime}})$ is of dimension $6\times 1$, which will be described shortly.
-To complete our model specification, we assume that the residual vector $\mathbf{u}_{t}$ are related to the structural shocks by
+To complete our model specification, we assume that the residual vector $\mathbf{u}_{t}$ is related to the structural shocks by
 $\mathbf{u}_{t}=\mathbf{B}\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{6\times 1},\mathbf{I}_{6}),$ (10)
@@
 It is known that the contemporaneous response matrix $\mathbf{B}$, hence, the structural shocks $\bm{\epsilon}_{t}$, cannot be separately identified without additional information.
-In this paper, we employ external instrument to identify monetary policy shock.
+In this paper, we employ an external instrument to identify monetary policy shocks.
 The next section provides the details of our identification approach.
 Our proposed model specified in equations (9)–(10) provides a unified framework for studying the macroeconomic stars and their corresponding cycle components.
-Furthermore, we allows the innovations of all trend and cycle components, i.e., $\mathbf{u}_{t}$, to be correlated.
-This contrasts with conventional studies on unobserved components models (e.g., Watson, , 1986; Stock and Watson, , 2007; Chan et al., , 2018; Laubach and Williams, , 2003; Zaman, , 2025), which typically assume these innovations to be independent.
+Furthermore, we allow the innovations of all trend and cycle components, i.e., $\mathbf{u}_{t}$, to be correlated.
+This contrasts with conventional studies on unobserved components models (e.g., watson1986univariate; stockwatson2; chan2018new; laubach2003measuring; zaman2025unified), which typically assume these innovations to be independent.
 Specifically, the contemporaneous response matrix $\mathbf{B}$ is assumed to be non-singular and unrestricted, which implies that $\bm{\Sigma}=\mathbf{B}\mathbf{B}^{\prime}$ is a full covariance matrix.
@@
 2.2 Identification via External Instrument
-In this section, we first discuss how the external instruments can be incorporated into our modeling framework for identification, and then describe our proposed SMUC-IV as an augmented strucutrual vector autoregression, a form which is used later for Bayesian estimation.
+In this section, we first discuss how the external instrument can be incorporated into our modelling framework for identification, and then describe our proposed SMUC-IV as an augmented structural vector autoregression, a form which is used later for Bayesian estimation.
 In this paper, we consider the case in which one external instrument is used to identify one structural shock, namely, the monetary policy shock of interest.
-For the case of using multiple external instruments to identify multiple structural shocks, we refer readers to Arias et al., (2021), Braun and Brüggemann, (2023), and Hou, (2024) for more details.
+For the case of using multiple external instruments to identify multiple structural shocks, we refer readers to arias2021inference, braun2023identification, and hou2024large for more details.
 To set the stage, we designate the last shock in $\bm{\epsilon}_{t}$, denoted by $\epsilon_{m,t}$, as the monetary policy shock of interest.
@@
 These two conditions are central to understanding how the external instrument can be used to identify the structural shock of interest $\epsilon_{m,t}$.
-Specifically, it conveys identifying information by distinguishing $\epsilon_{m,t}$ from the other shocks in $\bm{\epsilon}_{-m,t}$ through differences in their covariance structures with the the external instrument $m_{t}$.
+Specifically, it conveys identifying information by distinguishing $\epsilon_{m,t}$ from the other shocks in $\bm{\epsilon}_{-m,t}$ through differences in their covariance structures with the external instrument $m_{t}$.
 The relevance and exogeneity conditions together provide further information by imposing zero restrictions on the parameter vector $\bm{\gamma}$.
@@
 To summarize, our SMUC-IV is specified as equations (9), (10), and (11), subject to the zero restrictions given in (12).
-The external instrument equation (11) with the zero restrictions in (12) gives $m_{t}=\beta\epsilon_{m,t}+\alpha v_{t}$, which is consistent with the specification used in Caldara and Herbst, (2019).
+The external instrument equation (11) with the zero restrictions in (12) gives $m_{t}=\beta\epsilon_{m,t}+\alpha v_{t}$, which is consistent with the specification used in caldara2019monetary.
 Our proposed SMUC-IV can be represented more compactly as an augmented structural vector autoregression for $\widetilde{\bm{\eta}}_{t}=(\bm{\eta}_{t}^{\prime},m_{t})^{\prime}$:
@@
 This augmented structural vector autoregression representation is commonly used in Bayesian analysis to facilitate posterior inference.
-As indicated by recent studies (Arias et al., , 2021; Braun and Brüggemann, , 2023; Hou, , 2024), the use of external instruments can only achieve set-identification.
-For instance, in our case, it can be shown that the second to last column of $\widetilde{\mathbf{B}}$ that embeds the impact responses of monetary policy shocks, i.e., the last column of $\mathbf{B}$, which can only be identified up to sign changes.
+As indicated by recent studies (arias2021inference; braun2023identification; hou2024large), the use of external instruments can only achieve set-identification.
+For instance, in our case, it can be shown that the second-to-last column of $\widetilde{\mathbf{B}}$, which embeds the impact responses to the monetary policy shock and corresponds to the last column of $\mathbf{B}$, can only be identified up to a sign change.
 A common solution to this set-identification issue is to impose sign restrictions on the impulse responses, which are informed by economic theory.
 In this paper, rather than imposing dogmatic restrictions on response directions, we identify monetary policy shocks using an informative prior that accommodates estimation uncertainty.
-We will discuss this more in Section 3.1.
+We will discuss the prior in more detail in Section 3.1.
 3 Bayesian Estimation
@@
 This section describes the prior distributions assigned to the model parameters.
-Let $\bar{\bm{\tau}}_{0}=(g_{-1}^{*},g_{0}^{*},\pi_{0}^{*},r_{0}^{*})^{\prime}$ denote the vector of the initial state parameters.
-Let $\bm{\Phi}_{l,i,j}$ represent the $(i,j)$ element of the autoregressive coefficient matrix $\bm{\Phi}_{l}$ for $i,j=1,2,3$, $l=1,\ldots,p$, and let $\mathbf{B}_{i,j}$ denote the $(i,j)$ element of the contemporaneous impact matrix $\mathbf{B}$ for $i,j=1,\ldots,6$.
+Let $\bar{\bm{\tau}}_{0}=(g_{0}^{*},\pi_{0}^{*},r_{0}^{*})^{\prime}$ denote the vector of the initial state parameters, $\bm{\Phi}_{l,i,j}$ the $(i,j)$th element of the autoregressive coefficient matrix $\bm{\Phi}_{l}$ for $i,j=1,2,3$, $l=1,\ldots,p$, and $\mathbf{B}_{i,j}$ the $(i,j)$th element of the contemporaneous impact matrix $\mathbf{B}$ for $i,j=1,\ldots,6$.
 We assume the following independent priors:
@@
 The hyperparameters $\kappa_{1}$ and $\kappa_{2}$ control the shrinkage strength for the own-lag and cross-variable-lag coefficients, respectively.
-Empirical evidence from recent studies, such as Cross et al., (2020) and Chan, (2021), indicates that allowing the hyperparameters of shrinkage prior to be parameters to be estimated can substantially improves both forecasting accuracy and model fit.
+Empirical evidence from recent studies, such as cross2020macroeconomic and chan2021minnesota, indicates that allowing the hyperparameters of shrinkage prior to be parameters to be estimated can substantially improves both forecasting accuracy and model fit.
 Therefore, rather than fixing these hyperparameters at predetermined values, we treat $\kappa_{1}$ and $\kappa_{2}$ as unknown parameters and assign them the following prior distributions:
@@
 That means, the last column of $\mathbf{B}$ is the impact responses of the trend and cycle components to a monetary policy shock.
-Our prior reflects a belief that a one-standard-deviation increase in monetary policy shock $\epsilon_{m,t}$ results in a $100$ basis point increase in the cycle of interest rate $c_{r,t}$, while having no effect on $\Delta g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$, on impact.
-By the definition of $c_{r,t}$ in (3), the impact response of $c_{r,t}$ to a monetary policy shock is entirely attributable to the increase in $r_{t}$.
-This identifying information from our prior is similar to conventional studies on identifying monetary policy shock by normalize the magnitude of the interest rate response to a positive value.
-[^4] For instance, Bauer and Swanson, (2023) normalize the interest rate response to a monetary policy shock by $25$ basis points, while Miranda-Agrippino and Ricco, (2021) use a normalization of $100$ basis points.
-Unlike these approaches, we do not fix the size of the response of interest rate to a monetary policy shock, instead, we impose a relatively tight prior that centers the response at $100$ basis points.
-We also highlight that since we use an informative prior that centers the impact responses of $\Delta g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$ to a monetary policy shock at zeros, any nonzero effects found in our empirical study must be supported by the data.
+Our prior reflects a belief that a one-standard-deviation increase in monetary policy shock $\epsilon_{m,t}$ results in a $100$ basis point increase in the cycle of nominal interest rate $c_{i,t}$, while having no effect on $g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$, on impact.
+By the definition of $c_{i,t}$ in (3), the impact response of $c_{i,t}$ to a monetary policy shock is entirely attributable to the increase in $i_{t}$.
+This identifying information from our prior is similar to conventional studies on identifying monetary policy shock by normalize the magnitude of the nominal interest rate response to a positive value.
+[^4] For instance, bauer2023reassessment normalize the nominal interest rate response to a monetary policy shock by $25$ basis points, while miranda2021transmission use a normalization of $100$ basis points.
+Unlike these approaches, we do not fix the size of the response of nominal interest rate to a monetary policy shock, instead, we impose a relatively tight prior that centers the response at $100$ basis points.
+We also highlight that since we use an informative prior that centers the impact responses of $g_{t}^{*}$, $\pi_{t}^{*}$, $r_{t}^{*}$, $c_{g,t}$ and $c_{\pi,t}$ to a monetary policy shock at zeros, any nonzero effects found in our empirical study must be supported by the data.
 For the parameters in the external instrument equation, we set $V_{\beta}=1$, $V_{\alpha}=1$, $\alpha_{0}=0$, and $\beta_{0}=0.5\times\sigma_{m}$, where $\sigma_{m}$ denotes the standard deviation of the external instrument.
-This is comparable to the setting in Caldara and Herbst, (2019).
-For the initial state parameters $\bar{\bm{\tau}}_{0}$, we use an uninformative prior by setting $\mathbf{V}_{\bar{\tau}_{00}}=100\mathbf{I}_{4}$ and $\bar{\bm{\tau}}_{00}=(g_{1},g_{1},\pi_{1},r_{1})^{\prime}$ where $g_{1}$, $\pi_{1}$ and $r_{1}$ are the first observations of log real GDP, inflation, and the interest rate in the sample period, respectively.
+This is comparable to the setting in caldara2019monetary.
+For the initial state parameters $\bar{\bm{\tau}}_{0}$, we use an uninformative prior by setting $\mathbf{V}_{\bar{\tau}_{00}}=100\mathbf{I}_{3}$ and $\bar{\bm{\tau}}_{00}=(g_{1},\pi_{1},i_{1})^{\prime}$ where $g_{1}$, $\pi_{1}$ and $i_{1}$ are the first observations of log real GDP, inflation, and the nominal interest rate in the sample period, respectively.
 3.2 Posterior Sampler
 We now discuss the estimation of our model with the prior described in the previous section.
-To set the stage, let $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, where $\mathbf{y}_{t}=(g_{t},\pi_{t},r_{t},m_{t})^{\prime}$, denote the vector of observed data, $\bm{\Phi}=(\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})$ denote the collection of the autoregressive coefficient matrices and $\bm{\tau}=(\bar{\bm{\tau}}_{0}^{\prime},\bm{\tau}_{1}^{\prime},\ldots,\bm{\tau}_{T}^{\prime})^{\prime}$ denote a vector of state parameters.
+To set the stage, let $\mathbf{y}=(\mathbf{y}_{1}^{\prime},\ldots,\mathbf{y}_{T}^{\prime})^{\prime}$, where $\mathbf{y}_{t}=(g_{t},\pi_{t},i_{t},m_{t})^{\prime}$, denote the vector of observed data, $\bm{\Phi}=(\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})$ denote the collection of the autoregressive coefficient matrices and $\bm{\tau}=(\bar{\bm{\tau}}_{0}^{\prime},\bm{\tau}_{1}^{\prime},\ldots,\bm{\tau}_{T}^{\prime})^{\prime}$ denote a vector of state parameters.
 The joint posterior distribution can be simulated by sequentially sampling from the following conditional distributions:
@@
     $p(\kappa_{2}|\bm{\tau},\mathbf{B},\beta,\alpha,\bm{\Phi},\kappa_{1},\mathbf{y})$.
-In Step 1, we sample the state vector $\bm{\tau}$ using the precision sampling method (Chan and Jeliazkov, , 2009; McCausland et al., , 2011; Rue, , 2001) instead of traditional Kalman filter techniques.
-The precision-based sampling method has gain increasing popularity for estimating state space models due to its computational efficiency and straightforward implementation.
+In Step 1, we sample the state vector $\bm{\tau}$ using the precision sampling method (chan2009efficient; mccausland2011simulation; rue2001fast) instead of traditional Kalman filter techniques.
+The precision-based sampling method has gained increasing popularity for estimating state space models due to its computational efficiency and straightforward implementation.
 One complication under our framework, however, involves determining the conditional posterior distribution for $\bm{\tau}$.
-The conventional approach applies Bayes’ rule, which requires separately deriving both the conditional likelihood and the prior density of the state parameters.
-However, when the state and measurement equations are correlated, the derivation of the conditional posterior of the state parameters becomes more cumbersome (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Leiva-Leon and Uzeda, , 2023).
-In this paper, we introduce a novel and direct approach to derive the conditional posterior of $\bm{\tau}$ that circumvents the use of Bayes’ rule.
+The conventional approach requires separately deriving both the conditional likelihood and the prior density of the state parameters.
+However, when the state and measurement equations are correlated, the derivation of the conditional posterior of the state parameters becomes more cumbersome (grant2017bayesian; grant2017reconciling; leiva2023endogenous).
+In this paper, we introduce a novel and direct approach to derive the conditional posterior of $\bm{\tau}$.
 The key idea is to first obtain the joint conditional distribution of $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$, which will be shown later to be a Gaussian distribution, and then exploit the standard Gaussian conditioning properties to obtain the conditional posterior of $\bm{\tau}$.
@@
 where $\widetilde{\bm{\eta}}=(\widetilde{\bm{\eta}}_{1}^{\prime},\ldots,\widetilde{\bm{\eta}}_{T}^{\prime})^{\prime}$, $\widetilde{\bm{\Sigma}}=\widetilde{\mathbf{B}}\widetilde{\mathbf{B}}$,
-$\mathbf{H}_{1}=\begin{pmatrix}\mathbf{I}_{7}&\mathbf{0}_{7}&\mathbf{0}_{7}&\cdots&\cdots&\cdots&\cdots&\mathbf{0}_{7}\\ -\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{p}&\ddots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\vdots\\ \mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\mathbf{0}_{7}\\ \mathbf{0}_{7}&\cdots&\mathbf{0}_{7}&-\widetilde{\mathbf{A}}_{p}&\cdots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}\end{pmatrix},\quad\bm{\Xi}=\begin{pmatrix}\bm{\Xi}_{1}\\ \mathbf{0}_{4\times 4}\\ \bm{\Xi}_{2}\\ \mathbf{0}_{4\times 4}\\ \mathbf{0}_{7\times 4}\\ \vdots\\ \mathbf{0}_{7\times 4}\end{pmatrix}$
-with
-$\bm{\Xi}_{1}=\begin{pmatrix}-1&2&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix},\quad\bm{\Xi}_{2}=\begin{pmatrix}0&-1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}$.
+$\mathbf{H}_{1}=\begin{pmatrix}\mathbf{I}_{7}&\mathbf{0}_{7}&\mathbf{0}_{7}&\cdots&\cdots&\cdots&\cdots&\mathbf{0}_{7}\\ -\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ -\widetilde{\mathbf{A}}_{p}&\ddots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}&\mathbf{0}_{7}&\ddots&\vdots\\ \mathbf{0}_{7}&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\mathbf{0}_{7}\\ \mathbf{0}_{7}&\cdots&\mathbf{0}_{7}&-\widetilde{\mathbf{A}}_{p}&\cdots&-\widetilde{\mathbf{A}}_{2}&-\widetilde{\mathbf{A}}_{1}&\mathbf{I}_{7}\end{pmatrix},\quad\bm{\Xi}=\begin{pmatrix}\mathbf{I}_{3}\\ \mathbf{0}_{4\times 3}\\ \mathbf{0}_{7\times 3}\\ \vdots\\ \vdots\\ \mathbf{0}_{7\times 3}\end{pmatrix}$
 […]
@@
 where $\widehat{\bm{\tau}}=\bm{\mu}_{\tau}-\mathbf{K}_{\tau}^{-1}\mathbf{K}_{\tau,y}(\mathbf{y}-\bm{\mu}_{y})$.
-It can be check that the precision matrix $\mathbf{K}_{\tau}$ is a banded matrix.
-Therefore, one can efficiently sample from $\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1})$ using the precision-based approach of Chan and Jeliazkov, (2009).
+Note that the $(3T+3)\times(3T+3)$ precision matrix $\mathbf{K}_{\tau}$ is sparse and banded, as illustrated in Figure 1, which visualizes its structure with nonzero and zero elements shown in blue and white, respectively.
+[^5] Here the dimension of $\mathbf{K}_{\tau}$ is $432\times 432$, which matches the dimension used in our empirical application in Section 5.
+This banded structure enables efficient sampling from $\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1})$ using the precision-based approach of chan2009efficient.
+The procedure is as follows: first, we compute the lower triangular Cholesky factor $\mathbf{L}$ of $\mathbf{K}_{\tau}$ such that $\mathbf{K}_{\tau}=\mathbf{L}\mathbf{L}^{\prime}$, which is computationally efficient due to the banded structure of $\mathbf{K}_{\tau}$.
+A draw $\bm{\tau}\sim\mathcal{N}(\widehat{\bm{\tau}},\mathbf{K}_{\tau}^{-1})$ is then obtained by solving:
+$\bm{\tau}=\bm{\mu}_{\tau}-\mathbf{L}^{{}^{\prime}-1}\left(\mathbf{L}^{-1}\mathbf{K}_{\tau,y}(\mathbf{y}-\bm{\mu}_{y})+\mathbf{f}\right),\quad\mathbf{f}\sim\mathcal{N}\left(\mathbf{0}_{3(T+1)\times 1},\mathbf{I}_{3(T+1)}\right).$
+Since $\mathbf{L}$ is a banded lower triangular matrix, we can solve the required linear systems rapidly via forward and backward substitution, with $\mathcal{O}(T)$ complexity.
+This efficiently avoids directly computing $\mathbf{K}_{\tau}^{-1}$, which has $\mathcal{O}(T^{3})$ complexity.
+Figure 1: The sparsity pattern of the precision matrix $\mathbf{K}_{\tau}$.
 For sampling $(\mathbf{B},\beta,\alpha)$ in Step 2, we first note that these parameters correspond to the nonzero elements in the contemporaneous impact matrix $\widetilde{\mathbf{B}}$ of the augmented structural autoregression representation in (13).
 Consequently, sampling $(\mathbf{B},\beta,\alpha)$ is equivalent to sampling $\widetilde{\mathbf{B}}$ with the zero restrictions in (15) and the sign restriction on $\alpha$.
-Since these restrictions can be expressed as linear equality and inequality restrictions on $\widetilde{\mathbf{B}}$, we sample the nonzero parameters using the efficient sampler developed by Hou, (2024).
+Since these restrictions can be expressed as linear equality and inequality restrictions on $\widetilde{\mathbf{B}}$, we sample the nonzero parameters using the efficient sampler developed by hou2024large.
 We refer readers to this paper for more implementation details.
@@
 In this section, we present an approach for estimating the marginal likelihood of our proposed SMUC-IV and its restricted versions.
-Our method builds upon the modified harmonic mean estimator proposed by Gelfand and Dey, (1994), integrating it with the conditional Monte Carlo method of Chan, (2023) to enhance estimation accuracy.
+Our method builds upon the modified harmonic mean estimator proposed by gelfand1994bayesian, integrating it with the conditional Monte Carlo method of chan2023comparing to enhance estimation accuracy.
 We begin with an overview of the modified harmonic mean estimator and then outline our conditional Monte Carlo improved estimator for the SMUC-IV.
@@
 While the modified harmonic mean estimator described above is simulation-consistent and straightforward to implement, it might perform poorly when the model parameter $\bm{\theta}$ is high-dimensional.
-To improve estimation accuracy, we employ the conditional Monte Carlo method of Chan, (2023).
+To improve estimation accuracy, we employ the conditional Monte Carlo method of chan2023comparing.
 The key idea is to first analytically integrate out as many parameters in $\bm{\theta}$ as possible, and then construct a modified harmonic mean estimator using the resulting unconditional likelihood and prior on the remaining lower-dimensional parameters.
 4.2 Estimating the Marginal Likelihood for SMUC-IV
-We now outline our conditional Monte Carlo improved modified harmonic mean estimator for the marginal likelihood of the SMUC-IV. In our modeling setting, the marginal likelihood is given by
+We now outline our conditional Monte Carlo improved modified harmonic mean estimator for the marginal likelihood of the SMUC-IV. In our modelling setting, the marginal likelihood is given by
 $p(\mathbf{y})$ $=\int p(\mathbf{y}|\bm{\tau},\bm{\Phi},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2})p(\bm{\tau},\bm{\Phi},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2})d(\bm{\tau},\bm{\Phi},\mathbf{B},\beta,\alpha,\kappa_{1},\kappa_{2})$
@@
 The second equality follows from the conditional independence structure of the prior distributions.
-In the last equality, we have integrated out the state parameters $\bm{\tau}$ and hyperparameters $(\kappa_{1},\kappa_{2})$.
+In the last equality, we have integrated out the state parameters $\bm{\tau}$ and the hyperparameters $(\kappa_{1},\kappa_{2})$.
 Specifically, we have:
@@
 $p(\bm{\Phi})=\int p(\bm{\Phi}|\kappa_{1},\kappa_{2})p(\kappa_{1})p(\kappa_{2})d(\kappa_{1},\kappa_{2}).$ (22)
-Note that the expression in (21) can be obtained directly from result given in (20), which shows that the joint conditional distribution of $(\mathbf{y},\bm{\tau})$ is Gaussian.
-Consequently, $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$ can be derived using standard properties of the Gaussian distribution.
-See Appendix B for more details.
-Using the analytical expressions for (21) and (22), we can estimate the marginal likelihood of our SMUC-IV with the following conditional Monte Carlo improved modified harmonic mean estimator:
+Note that the expression in (21) can be obtained directly from the result in (20) by using the property that the marginal distribution of joint Gaussian variates is Gaussian.
+To be specific, let $\bm{\Lambda}=\mathbf{K}^{-1}$ be the covariance matrix of the joint conditional distribution of $(\bm{\tau}^{\prime},\mathbf{y}^{\prime})^{\prime}$, and partition $\bm{\Lambda}$ according to the dimensions of $\bm{\tau}$ and $\mathbf{y}$ as
+$\bm{\Lambda}=\begin{pmatrix}\bm{\Lambda}_{\tau}&\bm{\Lambda}_{\tau,y}\\ \bm{\Lambda}_{\tau,y}^{\prime}&\bm{\Lambda}_{y}\end{pmatrix}$.
+A standard property of the Gaussian distribution implies that the marginal distribution of $\mathbf{y}$ (of dimension $4T\times 1$) is a Gaussian distribution with mean vector $\bm{\mu}_{y}$ and covariance matrix $\bm{\Lambda}_{y}$.
+This yields
+$p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)=(2\pi)^{-2T}|\bm{\Lambda}_{y}|^{-\frac{1}{2}}e^{-\frac{1}{2}\left(\mathbf{y}-\bm{\mu}_{y}\right)^{\prime}\bm{\Lambda}^{-1}_{y}\left(\mathbf{y}-\bm{\mu}_{y}\right)}.$
+For the analytical expression of the marginal prior $p(\bm{\Phi})$ in (22), we refer readers to Appendix B for details.
+Given the analytical expressions for (21) and (22), we can estimate the marginal likelihood of our SMUC-IV with the following conditional Monte Carlo improved modified harmonic mean estimator:
 $\widehat{p(\mathbf{y})}_{CMGD}=\left(\frac{1}{R}\sum_{i=1}^{R}\frac{q(\bm{\Phi}^{(i)},\mathbf{B}^{(i)},\alpha^{(i)},\beta^{(i)})}{p(\mathbf{y}|\bm{\Phi}^{(i)},\mathbf{B}^{(i)},\alpha^{(i)},\beta^{(i)})p(\mathbf{B}^{(i)})p(\bm{\Phi}^{(i)})p(\alpha^{(i)})p(\beta^{(i)})}\right)^{-1},$ (23)
@@
 Note that the key difference between this estimator and the standard modified harmonic mean estimator is the use of the likelihood function $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$, which is unconditional on the high-dimensional state parameters $\bm{\tau}$, and the marginal prior $p(\bm{\Phi})$.
-This formulation substantially reduces the dimensionality of the numerical integration, which in turn reduces the Monte Carlo variance of the estimator, resulting in greater numerical stability and estimation precision (see Chan, (2023) for more discussion).
+This formulation substantially reduces the dimensionality of the numerical integration, which in turn reduces the Monte Carlo variance of the estimator, resulting in greater numerical stability and estimation precision (see chan2023comparing for more discussion).
 We now turn to the choice of the tuning density function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$, a critical component of our estimator (23) that is essential for the accuracy of marginal likelihood estimation.
 Although the posterior density is the theoretically optimal choice, it is computationally intractable.
-Therefore, we follow Chan, (2023) and approximate the posterior using a truncated Gaussian density.
+Therefore, we follow chan2023comparing and approximate the posterior using a truncated Gaussian density.
 Specifically, we consider a tuning density function that takes the following form:
@@
 where each $q_{j}(\cdot)$ for $j\in\{\Phi,B,\alpha,\beta\}$ is a truncated Gaussian density with mean and covariance matrix set to the estimated posterior mean and covariance matrix.
-[^5] Using an appropriately truncated tuning density can ensure that the modified harmonic mean estimator has a finite variance (Geweke, , 1999).
+[^6] Using an appropriately truncated tuning density can ensure that the modified harmonic mean estimator has a finite variance (geweke1999using).
 The full details for $q_{\Phi}(\bm{\Phi})$, $q_{B}(\mathbf{B})$, $q_{\alpha}(\alpha)$, and $q_{\beta}(\beta)$ are provided in Appendix B.
@@
 In this section, we first describe the data in Section 5.1 and then conduct a model comparison exercise in Section 5.2 to validate our proposed SMUC-IV framework.
-Section 5.3 presents our estimates of the macroeconomic stars—namely, the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate.
-Section 5.4 investigates the effects of monetary policy shocks on these stars.
-Subsequently, Section 5.5 assesses the role of the monetary policy shocks in driving the evolution of the macroeconomic stars.
+In particular, this exercise provides strong evidence in favor of the influence of monetary policy shocks on the stars.
+Section 5.3 presents our estimates of the macroeconomic stars—namely, the potential GDP growth, trend inflation, and the neutral interest rate.
+Section 5.4 examines the effects of monetary policy shocks on these stars.
+Subsequently, Section 5.5 assesses the role of the monetary policy in driving the evolution of the macroeconomic stars.
+Section 5.6 provides a discussion of the results and their policy implications.
 Finally, robustness analyses are reported in Section 5.7.
 5.1 Data and External Instrument
-We estimate the empirical model using the following quarterly data from 1987:Q4 to 2023:Q4: real GDP, the GDP deflator, the federal funds effective rate, the shadow interest rate, and the orthogonal monetary policy surprise (Bauer and Swanson, , 2023).
-To capture both conventional and unconventional monetary policy when the nominal federal funds rate is at the effective lower bound (ELB), we use the Wu and Xia, (2016) shadow short rate.
-Consistent with the quarterly frequency of our data we use four lags in our model.
-We transform real GDP as $100\times\log(x_{t})$ and compute the quarterly GDP deflator as annualized log growth, $400\times\log(x_{t}/x_{t-1})$.
+We estimate the empirical model using the following quarterly data from 1988:Q1 to 2023:Q3: real GDP, the GDP deflator, the federal funds effective rate, the shadow interest rate, and the orthogonal monetary policy surprise (bauer2023reassessment).
+To capture both conventional and unconventional monetary policy when the nominal federal funds rate is at the effective lower bound (ELB), we use the wu2016measuring shadow short rate.
+Consistent with the quarterly frequency of our data, we use four lags in our model.
+We transform real GDP as $100\log(x_{t})$ and compute real GDP growth as the first difference of this series.
+The quarterly GDP deflator inflation rate is measured as annualized log growth, 400 $\log(x_{t}/x_{t-1})$.
 The federal funds effective rate, shadow interest rate, and orthogonal monetary policy surprise are originally monthly series; we obtain their quarterly versions by averaging the three monthly observations within each quarter.
 Real GDP, the GDP deflator, and the federal funds effective rate are available from the FRED database.
-We consider different measures of inflation and the interest rate as well as consider the bond premium (Favara et al., , 2016) as a control variable in the vector autoregression, see section 5.6.
+We consider different measures of inflation and the interest rate as well as consider the bond premium (favara2016updating) as a control variable in the vector autoregression, see section 5.7.
 The shadow interest rate is from the website of the Federal Reserve Bank of Atlanta, the excess bond premium is from the website of the Board of Governors of the Federal Reserve System, and the orthogonal monetary policy surprise series is from the website of the Federal Reserve Bank of San Francisco.
-In this paper, we use the orthogonal monetary policy surprise series developed by Bauer and Swanson, (2023) as our external instrument.
+In this paper, we use the orthogonal monetary policy surprise series developed by bauer2023reassessment as our external instrument.
 This series identifies monetary policy shocks from high-frequency asset price movements in narrow windows around policy announcements.
-High-frequency interest rate changes around FOMC announcements are a common tool for identifying monetary policy effects, but recent studies have questioned their exogeneity and relevance as instruments, particularly for estimating macroeconomic impacts (e.g., Ramey, , 2016; Miranda-Agrippino and Ricco, , 2021).
+High-frequency interest rate changes around FOMC announcements are a common tool for identifying monetary policy effects, but recent studies have questioned their exogeneity and relevance as instruments, particularly for estimating macroeconomic impacts (e.g., ramey2016macroeconomic; miranda2021transmission).
 For instance, monetary policy surprises may be correlated with publicly available macroeconomic and financial data released before FOMC announcements.
-To address these concerns, Bauer and Swanson, (2023) (i) expand the set of monetary policy events to include speeches by the Federal Reserve Chair—roughly doubling the number of announcements—and (ii) orthogonalize the resulting surprises with respect to pre-announcement macroeconomic and financial data to account for predictability via the “Fed response to news” channel.
+To address these concerns, bauer2023reassessment (i) expand the set of monetary policy events to include speeches by the Federal Reserve Chair—roughly doubling the number of announcements—and (ii) orthogonalize the resulting surprises with respect to pre-announcement macroeconomic and financial data to account for predictability via the “Fed response to news” channel.
+In Section 5.7, we also conduct a robustness analysis by considering an alternative instrument in our application, and we find that our main empirical results remain robust under this alternative choice of instrument.
 5.2 Model Comparison Exercise
@@
 We refer to this model as SMUC-IV-R1.
-Next, we consider a restricted version of our SMUC-IV that imposes zero correlation between the external instrument and all structural shocks by setting $\beta=0$.
+Next, we consider a restricted version of our SMUC-IV model that imposes zero correlation between the external instrument and all structural shocks by setting $\beta=0$; we refer to this specification as SMUC-IV-R2.
 A comparison between our proposed SMUC-IV and SMUC-IV-R2 serves to test the relevance condition of the instrument.
@@
 A few findings are also worth highlighting.
-First, by comparing the SMUC-IV with the SMUC-IV-R1, the difference in the log marginal likelihoods is about $13$, strongly supporting the non-zero contemporaneous effects of monetary policy shocks on the trend components.
+First, by comparing the SMUC-IV with the SMUC-IV-R1, the difference in the log marginal likelihoods is about $18$, strongly supporting the non-zero contemporaneous effects of monetary policy shocks on the trend components.
 Second, the SMUC-IV-R2 performs better than the SMUC-IV-R3 and SMUC-IV-R4.
-This suggests that allowing for correlated innovations among all trend and cycle components is an important modeling feature supported by the data.
-In particular, the log marginal likelihood of SMUC-IV-R2 is about $24$ higher than that of SMUC-IV-R3, indicating that the assumption of independent innovations for the trend and cycle components is empirically implausible in our empirical analysis.
-Lastly, the SMUC-IV outperforms SMUC-IV-R1, suggesting that the relevance condition of the instrument is satisfied, although the difference in the log marginal likelihoods between these two models is only about 33.666In Section 5.6, we also conduct a robustness analysis by considering an alternative instrument in our application, and we find that our main empirical results remain robust under this alternative choice of instrument.
+This suggests that allowing for correlated innovations among all trend and cycle components is an important modelling feature supported by the data.
+In particular, the log marginal likelihood of SMUC-IV-R2 is about $22$ higher than that of SMUC-IV-R4, indicating that the assumption of independent innovations for the trend and cycle components is empirically implausible in our empirical analysis.
+Lastly, the SMUC-IV outperforms SMUC-IV-R2, suggesting that the relevance condition of the instrument is satisfied.
 Table 2: Estimated log marginal likelihoods
 SMUC-IV SMUC-IV-R1 SMUC-IV-R2 SMUC-IV-R3 SMUC-IV-R4
--207 -220 -210 -223 -231
+-205 -223 -208 -225 -230
 5.3 Estimates of Macroeconomic Stars
@@
 Accordingly, we examine the persistent evolution of macroeconomic stars through the lens of trend dynamics.
-Using US data, we jointly estimate four stars: the level of potential GDP, $g^{*}$, the growth rate of potential GDP, $\Delta g^{*}$, and trend inflation, $\pi^{*}$, and the neutral interest rate, $r^{*}$.
-These stars are measured as long-run trends in the spirit of the Beveridge-Nelson decomposition (Beveridge and Nelson, , 1981).
+Using US data, we jointly estimate three stars: the potential GDP growth, $g^{*}$, the trend inflation, $\pi^{*}$, and the neutral interest rate, $r^{*}$.
 […]
-Panel (a) of Figure 1 presents the estimates of $^{*}.
-gg^{*}$ rises over time but experiences two notable drops during the 2007–2009 Global Financial Crisis and the COVID-19 pandemic.
-In both cases, the level of $g^{*}$ fails to return to its pre-crisis trajectory, suggesting that transitory shocks may translate into permanent reductions in the long-run growth path.
-Panel (b) of Figure 1 shows estimates of $\Delta g^{*}$.
-From the early to late 1990s, $\Delta g^{*}$ rises noticeably, likely reflecting the internet technology boom.
+Figure 2: The plot compares the actual data with trend estimates along 68% credible bands.
+Panel (a) of Figure 2 shows estimates of $g^{*}$.
+From the early to late 1990s, $g^{*}$ rises noticeably, likely reflecting the internet technology boom.
 After the 1990s, however, it follows a downward trend until around 2010.
 Then it follows an upward trend until the early 2020s.
-Our estimates show a similar tendency to those in (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Maffei-Faccioli, , 2025), albeit with considerable fluctuations.
+Our estimates show a similar tendency to those in (grant2017bayesian; grant2017reconciling; maffei2025identifying).
 Sharp declines are evident during major crises such as the dot-com bust in the early 2000s, the Global Financial Crisis of 2007–2009, and the COVID-19 pandemic.
-These episodes coincide with well-documented hysteresis effects, where severe recessions leave lasting scars on the economy’s productive capacity (Cerra and Saxena, , 2005, 2008; Cerra et al., , 2023).
-Panel (c) of Figure 1 reports the estimates of $\pi^{*}$.
-Consistent with prior work (e.g., Stock and Watson, , 2007, 2016; Chan et al., , 2018; Eo et al., , 2023), $\pi^{*}$ displays remarkable stability: after falling in the 1990s, it remains anchored near 2% from the 2000s until the pandemic period.
+These episodes coincide with well-documented hysteresis effects, where severe recessions leave lasting scars on the economy’s productive capacity (cerra2005did; cerra2008growth; cerra2023hysteresis).
+Panel (b) of Figure 2 reports the estimates of $\pi^{*}$.
+Consistent with prior work (e.g., stockwatson2; stock2016core; chan2018new; eo2023understanding), $\pi^{*}$ displays remarkable stability: after falling in the 1990s, it remains anchored near 2% from the 2000s until the pandemic period.
 This pattern highlights the Federal Reserve’s success in stabilizing long-run inflation expectations.
-Finally, Panel (d) of Figure 1 presents estimates of $r^{*}$.
+Finally, Panel (c) of Figure 2 presents estimates of $r^{*}$.
 The estimates show a steady decline from the 1990s onward, followed by relative stability during the 2010s-2020s.
-Our estimates of $r^{*}$ capture the secular decline in the neutral interest rate documented in the literature on estimating $r^{*}$ (e.g., Laubach and Williams, , 2003; Lubik and Matthes, , 2015; Del Negro et al., , 2017; Morley et al., , 2024).
-Figure 1: The plot compares the actual data with trend estimates along 68% credible bands.
-5.4 Do Monetary Policy Shocks Affect Macroeconomic Stars?
-Having obtained plausible estimates of the macroeconomic stars, we now turn to the paper’s central aim: exploring the effects of monetary policy shocks on these stars.
-Specifically, we examine how monetary policy influences $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$.
-Monetary policy shocks are identified using an external instrument based on high-frequency policy surprises developed by Bauer and Swanson, (2023).
+Our estimates of $r^{*}$ capture the secular decline in the neutral interest rate documented in the literature on estimating $r^{*}$ (e.g., laubach2003measuring; lubik2015calculating; del2017safety; morley2024simple).
+We also report the estimated cyclical components in Figure 3.
+Overall, the cycles appear stationary, suggesting that our proposed model effectively decomposes the time series of interest into trends and cycles.
+The cyclical components capture major economic fluctuations, including the recessions in the early 1990s, the global financial crisis of 2008-2009, and the sharp contraction during the COVID-19 pandemic in 2020.
+Figure 3: The plot shows cycle estimates with 68% credible bands.
+5.4 The Effects of Monetary Policy Shocks on the Macroeconomic Stars
+Having obtained plausible estimates of the macroeconomic stars, we now examine the effects of monetary policy shocks on $g^{*}$, $\pi^{*}$, and $r^{*}$.
+We identify monetary policy shocks using an external instrument based on the high-frequency policy surprises constructed by bauer2023reassessment.
 Figure 4: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. Note that the responses are constant over time.
-Figure 2 shows the posterior distributions of the impulse responses of $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$ to a one-standard-deviation contractionary monetary policy shock.
-This shock raises the median response of the interest rate by about 0.29 basis points on impact (see Figure 3).
+Figure 4 shows the posterior distributions of the impulse responses of $g^{*}$, $\pi^{*}$, and $r^{*}$ to a one-standard-deviation contractionary monetary policy shock.
+This shock raises the median response of the nominal interest rate by about 29 basis points on impact (see Figure 5(f)).
 By construction, the responses are constant over time, as the model assumes the stars follow a random walk.
-Panel (a) of Figure 2 shows that $\Delta g^{*}$ falls by about 0.1 percentage point.
-The negative effects on the growth rate of potential output align with the main finding of Jordà et al., (2024) that monetary policy can have long-run effects on the real economy, suggesting an innovation channel whereby monetary tightening curbs spendings on R&D and innovation, limits productivity gains, and reduces productive investment, thereby depressing the growth rate of potential output (see, e.g., Stadler, , 1990; Moran and Queralto, , 2018; Ma and Zimmermann, , 2023; Fornaro and Wolf, , 2023; Meier and Reinelt, , 2024).
+Panel (a) of Figure 4 shows that $g^{*}$ falls by about 0.05 percentage points.
+The negative effects on potential GDP growth align with the empirical evidence in moran2018innovation, garga2021output, ma2023monetary, jorda2024long, and meier2024monetary, which suggests that tighter monetary policy restrains innovation investment and, in turn, productivity.
 Panel (b) of Figure 4 shows that contractionary monetary policy shocks also reduce $\pi^{*}$ by about 0.2 percentage points.
-At first glance, significant effects of monetary policy shocks on trend inflation may appear undesirable, as they seem to imply de-anchoring of long-run inflation expectations.
-However, when long-run expectations are persistently off target, such effects may reflect a re-anchoring channel, in which policy deliberately shifts expectations back toward the target.
-As shown by our counterfactual analysis (see Figure 4), monetary policy shocks pull trend inflation down make it move around the 2% target from mid-1990s to the onset of Covid-19 pandemic.
-This finding is consistent with evidence that such shocks lower survey-based long-run inflation expectations, pointing to a re-anchoring channel in which monetary policy shifts long-run inflation expectations back toward target when they drift persistently away from it (see, e.g., Jarociński and Karadi, , 2020; Diegel and Nautz, , 2021).
-Panel (c) shows that $r^{*}$ falls by about 0.3 basis points.
-This decline, caused by contractionary monetary policy shocks, may operate through two complementary mechanisms.
-First, the innovation channel in panel (a) indicates that contractionary monetary policy shocks lowers potential output growth; because trend growth is a key determinant of the neutral rate (e.g., Laubach and Williams, , 2003), this directly pushes $r^{*}$ down.
-Second, the re-anchoring channel in panel (b) indicates that these shocks lower $\pi^{*}$, thereby raising the real interest rate and also encouraging firms to cut spending on R&D and productive investment —reinforcing the innovation channel and further depressing $r^{*}$.
-A more practical question concerns how actual macroeconomic variables respond to contractionary monetary policy shocks once policy-induced shifts in the macroeconomic stars are taken into account.
-Figure 3 shows that the impulse responses of the actual variables align with standard macroeconomic predictions: the nominal interest rate rises, output falls, and inflation declines.
-Panel (a) of Figure 3 shows that output falls by about 1.5 percent after five years.
-Panel (b) indicates that inflation declines by roughly 0.17 percentage points after five years, following the 0.29-basis-point increase in the nominal interest rate shown on impact in Panel (c).
-Notably, when monetary policy shocks are allowed to affect the macroeconomic stars, the responses of the actual variables become markedly more persistent.
-Figure 3: The plots shows the impulse responses with 68% impulse responses of the actual variable to a monetary policy shock.
-5.5 Are the Monetary Policy Shocks a Key Driver for Shaping the Macroeconomic Stars?
-Given the meaningful effects of monetary policy shocks on the macroeconomic stars, an important question is whether these shocks have played a important role in shaping the historical dynamics of the stars.
+The decline in trend inflation is consistent with evidence that contractionary monetary policy shocks reduce long-term inflation expectations (e.g., jarocinski2020deconstructing; diegel2021long).
+Panel (c) shows that $r^{*}$ falls by about 25 basis points.
+The decline in the neutral interest rate after contractionary monetary policy shocks could be interpreted as reflecting the adverse effects of monetary tightening on innovation investment.
+As a complement to our analysis of the effects of monetary policy shocks on the stars, we also report the cyclical responses and the responses of the observed variables to contractionary monetary policy shocks in Figure 5.
+Figure 5: The plots show the impulse responses of the cyclical components and the observed variables to a monetary policy shock, along with 68% credible intervals.
+Figure 5 shows that the impulse responses align with standard macroeconomic predictions: the nominal interest rate rises, real GDP growth falls, and inflation declines.
+The figure also indicates that contractionary monetary policy shocks have more persistent effects on the business cycle than expected, taking about 3–5 years to fade out.
+Moreover, the responses of the observed variables appear to be largely permanent.
+Overall, our results suggest that monetary tightening may generate medium-run adverse effects on aggregate demand, which in turn suppress investment in innovation and weaken long-run productivity, leaving potentially permanent scars on the economy.
+5.5 The Role of Monetary Policy Shocks in Shaping Macroeconomic Stars
+Given the meaningful effects of monetary policy shocks on the macroeconomic stars, an important question is whether these shocks have played an important role in shaping the historical dynamics of the stars.
 To address this, we conduct a counterfactual analysis based on the historical decomposition of the trend components, as shown in Figure 6.
 Figure 6: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
-Figure 4 compares the actual trend paths with counterfactual paths, which are generated by shutting down all identified monetary policy shocks, following the methodology outlined in Kilian and Lütkepohl, (2017).
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
+Figure 6 compares the actual trend paths with counterfactual paths, which are generated by shutting down all identified monetary policy shocks, following the methodology outlined in kilian2017structural.
 The shaded regions represent 68% credible bands around the counterfactual estimates.
 Our counterfactual analysis quantifies the contributions of monetary policy shocks over time.
-For instance, in Figure 4, Panel (a) shows that, absent these shocks, $\Delta g^{*}$ would have been over 1 percent higher during the crisis; Panel (b) indicates that $g^{*}$ would have been persistently higher, with a cumulative gap of about 0.5 percent over the sample period.
-Panel (c) shows that $\pi^{*}$, which remained below its counterfactual path since the early 1990s, would have hovered closer to 3 percent in the absence of repeated monetary tightening.
-Panel (d) shows that $r^{*}$ would have been more than 2 basis points higher during the financial crisis.
-Taken together, the counterfactual analysis emphasizes that monetary policy is a main driver of the macroeconomic stars and tends to play a dual role: it effectively stabilizes long-run inflation expectations but can also impose a persistent drag on the real economy’s long-run trends when over-tightening occurs.
-Policymakers should therefore weigh the benefits of anchoring inflation against the potential costs for the real economy in the long run.
+For instance, in Figure 6, Panel (a) shows that, absent these shocks, $g^{*}$ would have been over 0.8 percentage points higher during the crisis.
+Panel (b) shows that $\pi^{*}$, which has remained below its counterfactual path since the early 1990s, would have hovered closer to 3.5 percent in the absence of monetary tightening.
+Panel (c) shows that $r^{*}$ would have been around 3 percentage points higher during the financial crisis.
+5.6 Results Discussion and Policy Implications
+The negative effects of contractionary monetary policy shocks on $g^{*}$ and $r^{*}$ shown in Figure 4 can be rationalized by the mechanism emphasized in the Keynesian endogenous growth model with nominal rigidities in fornaro2023scars.
+In that setting, long-run growth is endogenous and can be influenced by both demand- and supply-side forces.
+With nominal rigidities, output can deviate from potential, so monetary policy has real effects.
+In their framework, monetary tightening can trigger a “supply–demand doom loop”: a contraction reduces aggregate demand, compressing firms’ profits and weakening incentives to invest.
+The resulting decline in investment lowers expected productivity growth and household wealth, which further depresses demand and induces additional cutbacks in investment.
+As a result, the $g^{*}$ and $r^{*}$ fall.
+Moreover, by amplifying the effects of adverse disturbances on the real economy, this vicious spiral helps interpret panel (a) and (c) in Figure 6: while the actual paths of $g^{*}$ and $r^{*}$ (with contractionary monetary policy shocks) and the counterfactual paths (excluding the shocks’ effects) share a similar downward tendency, the actual series remain persistently below their counterfactual counterparts.
+Turning to trend inflation, at first glance, significant negative effects of monetary policy shocks in panel (b) of Figure 4 may appear undesirable, as they could be read as evidence of de-anchoring in long-run inflation expectations.
+However, if long-run expectations are persistently off target, such effects may instead reflect a re-anchoring channel, whereby policy deliberately shifts expectations back toward the target (for detailed discussion, see, e.g., diegel2021long).
+Consistent with this interpretation, our counterfactual analysis (panel (b) of Figure 6) indicates that monetary policy shocks pulled trend inflation down and kept it closer to the 2% target from the mid-1990s through the onset of the COVID-19 pandemic.
+Taken together, our results suggest that central banks face a dilemma: bringing inflation back to target may come at the cost of lower potential GDP growth and a lower neutral interest rate.
+This, in turn, points to a potential role for fiscal interventions that support business investment and the economy’s productive capacity during disinflation episodes.
 5.7 Robustness Analysis
 We assess the robustness of our baseline findings by re-estimating the model under a range of alternative specifications.
-Figures C.1-C.8 show results of the IRFs of the trends and C.9-C.16 show the historical decomposition of all alternative specifications.
+Figures C.1-C.8 show results of the IRFs of the trends and Figures C.9-C.16 show the historical decomposition of all alternative specifications.
 Across all variations, the qualitative responses of the macroeconomic stars to a contractionary monetary policy shock remain similar to those in the baseline, indicating that our main results are not driven by specific modelling choices.
 First, estimating the shrinkage parameter $V_{b}$ via a hierarchical Bayes approach yields posterior impulse responses that closely match the baseline, with slightly wider credible intervals (Figure C.1) and very similar historical decompositions (Figure C.9).
-Second, restricting the sample to pre-COVID observations ending in 2019Q4 produces similar negative responses of $\Delta g^{*}$, $\pi^{*}$, and $r^{*}$ (Figure C.2) as well as an similar counterfactual path of the trends (Figure C.10), suggesting that the pandemic period does not drive the results.
-Third, increasing the VAR lag length from four to eight (Figure C.3 and igure C.11) does not materially alter the estimated responses or historical decompositions.
-Fourth, replacing the federal funds rate with the one-year or two-year Treasury yield (Figures C.4, C.5, C.12 and C.13) preserves the qualitative pattern of the responses and the pattern of the counterfactual pathes.
+Second, restricting the sample to pre-COVID observations ending in 2019Q4 produces similar negative responses of $g^{*}$, $\pi^{*}$, and $r^{*}$ (Figure C.2) as well as a similar counterfactual path of the trends (Figure C.10), suggesting that the pandemic period does not drive the results.
+Third, increasing the VAR lag length from four to eight (Figures C.3 and C.11) does not materially alter the estimated responses or historical decompositions.
+Fourth, replacing the federal funds rate with the one-year or two-year Treasury yield (Figures C.4, C.5, C.12 and C.13) preserves the qualitative pattern of the responses and the pattern of the counterfactual paths.
 Fifth, substituting the GDP deflator with the PCE price index (Figures C.6 and C.14) yields comparable estimates.
 Sixth, using the unadjusted monetary policy surprise measure from Bauer and Swanson (2023) instead of the orthogonalized version (Figures C.7 and C.15) results in nearly identical posterior distributions.
-Finally, controlling for financial conditions by adding the excess bond premium (Favara et al., , 2016) to the cycle part of the model confirms our main empirical findings( Figures C.8 and C.16).
-Overall, these robustness analyses confirm that the estimated effects of monetary policy shocks on the macroeconomic stars are stable across alternative sample periods, lag specifications, interest rate measures, inflation measures, external instruments, and controlling for financial conditions using the excess bond premium.
+Finally, controlling for financial conditions by adding the excess bond premium (favara2016updating) to the cycle part of the model confirms our main empirical findings( Figures C.8 and C.16).
+Overall, these robustness analyses confirm that the estimated effects of monetary policy shocks on the macroeconomic stars are stable across alternative sample periods, lag specifications, nominal interest rate measures, inflation measures, external instruments, and controlling for financial conditions using the excess bond premium.
 6 Conclusions
 This paper develops SMUC-IV to explore the effects of monetary policy shocks on the macroeconomic stars, measured as long-run trends, in a unified framework.
-Using our SMUC-IV, we show that monetary tightening can have negative effects on the growth rate of potential output, trend inflation, and the neutral interest rate, and that monetary policy shocks are an important driver of the growth rate of potential output (along with the level of potential output), trend inflation, and the neutral interest rate.
-From a policy perspective, the results indicate that policymakers can use monetary policy to re-anchor long-run inflation expectations at target, but also caution that too much tightening could harm the real economy in the long run.
-Recently, growing attention has been paid to the long-run effects not only of monetary policy shocks but also of other structural shocks, such as financial, fiscal, and demand shocks (Cerra and Saxena, , 2005, 2008; Antolin-Diaz and Surico, , 2025; Furlanetto et al., , 2025).
+Using our SMUC-IV, we show that monetary tightening can have negative effects on the potential GDP growth, trend inflation, and the neutral interest rate, and that monetary policy is an important driver of these stars.
+From a policy perspective, our results indicate that policymakers can bring elevated inflation back to target through monetary tightening, but potentially at the cost of long-run damage to the real economy.
+This dilemma points to a role for complementary fiscal interventions that support business investment and the economy’s productive capacity during disinflation episodes.
+Recently, growing attention has been paid to the long-run effects not only of monetary policy shocks but also of other structural shocks, such as financial, fiscal, and demand shocks (cerra2005did; cerra2008growth; antolin2025long; furlanetto2025estimating).
 Against this backdrop, future research could extend our framework to examine how different types of structural shocks shape long-run macroeconomic equilibria.
@@
 […]
-where $\bar{\bm{\eta}}_{t}=(\Delta g_{t}^{*}-\Delta g_{t-1}^{*},\pi_{t}^{*}-\pi_{t-1}^{*},r_{t}^{*}-r_{t-1}^{*},\mathbf{c}_{t}^{\prime},m_{t})^{\prime}$, $\widetilde{\bm{\Sigma}}=\widetilde{\mathbf{B}}\widetilde{\mathbf{B}}^{\prime}$ and $\bar{\mathbf{A}}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i},0)$ for $i=1,\ldots,p$.
+where $\bar{\bm{\eta}}_{t}=(g_{t}^{*}-g_{t-1}^{*},\pi_{t}^{*}-\pi_{t-1}^{*},r_{t}^{*}-r_{t-1}^{*},\mathbf{c}_{t}^{\prime},m_{t})^{\prime}$, $\widetilde{\bm{\Sigma}}=\widetilde{\mathbf{B}}\widetilde{\mathbf{B}}^{\prime}$ and $\bar{\mathbf{A}}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i},0)$ for $i=1,\ldots,p$.
 Stacking equation (25) over $t=1,\ldots,T$, we have
@@
 where $\bar{\bm{\eta}}=(\bar{\bm{\eta}}_{1}^{\prime},\ldots,\bar{\bm{\eta}}_{T}^{\prime})^{\prime}$, $\bar{\mathbf{a}}=\text{vec}\left((\bar{\mathbf{A}}_{1},\ldots,\bar{\mathbf{A}}_{p})^{\prime}\right)$, $\bar{\mathbf{X}}=(\bar{\mathbf{X}}_{1}^{\prime},\ldots,\bar{\mathbf{X}}_{T}^{\prime})^{\prime}$ with $\bar{\mathbf{X}}_{t}=\mathbf{I}_{7}\otimes(\bar{\bm{\eta}}_{t-1}^{\prime},\ldots,\bar{\bm{\eta}}_{t-p}^{\prime})$.
-Let $\bm{\phi}=\text{vec}\left((\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})^{\prime}\right)$ and $\mathbf{S}_{\phi}$ be the selection matrix such that $\bar{\mathbf{a}}=\mathbf{S}_{\phi}\bm{\phi}$.
-Then we can write (26) as
+Let $\bm{\phi}=\text{vec}\left((\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p})^{\prime}\right)$ and $\mathbf{S}_{\phi}$, then there is a selection matrix $\mathbf{S}_{\phi}$ such that $\bar{\mathbf{a}}=\mathbf{S}_{\phi}\bm{\phi}$.
+It follows that we can write (26) as
 […]
@@
 Appendix Appendix B Details of Marginal Likelihood Estimation
-In this Appendix, we first present the analytical expressions for (21) and (22) and then detail the construction of the tuning function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$.
-Expressions for $p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)$ and $p(\bm{\Phi})$
-From result in (20), let $\bm{\Lambda}=\mathbf{K}^{-1}$ and partition $\bm{\Lambda}$ according to the dimensions of $\bm{\tau}$ and $\mathbf{y}$ as
-$\bm{\Lambda}=\begin{pmatrix}\bm{\Lambda}_{\tau}&\bm{\Lambda}_{\tau,y}\\ \bm{\Lambda}_{\tau,y}^{\prime}&\bm{\Lambda}_{y}\end{pmatrix}$.
-The property of Gaussian distribution implies that the marginal distribution of $\mathbf{y}$ (of dimension $4T\times 1$) is a Gaussian with means $\bm{\mu}_{y}$ and covariance matrix $\bm{\Lambda}$.
-This yields
-$p(\mathbf{y}|\bm{\Phi},\mathbf{B},\alpha,\beta)=(2\pi)^{-2T}|\bm{\Lambda}_{y}|^{-\frac{1}{2}}e^{-\frac{1}{2}\left(\mathbf{y}-\bm{\mu}_{y}\right)^{\prime}\bm{\Lambda}^{-1}_{y}\left(\mathbf{y}-\bm{\mu}_{y}\right)}.$
-Next, we derive the expression for the marginal prior $p(\bm{\Phi})$. Given the priors of $\bm{\Phi}$ and $(\kappa_{1},\kappa_{2})$ described in Section 3.1, we have
+In this Appendix, we first present the analytical expression of the marginal prior $p(\bm{\Phi})$ in (22), and then detail the construction of the tuning function $q(\bm{\Phi},\mathbf{B},\alpha,\beta)$.
+Expressions for $p(\bm{\Phi})$
+Given the priors of $\bm{\Phi}$ and $(\kappa_{1},\kappa_{2})$ described in Section 3.1, we have
 $p(\bm{\Phi})$ $=\int p(\bm{\Phi}|\kappa_{1},\kappa_{2})p(\kappa_{1})p(\kappa_{2})d(\kappa_{1},\kappa_{2})$
@@
 $\hat{\nu}_{1}=\frac{3p}{2}-1,\qquad\hat{S}_{1}=\frac{1}{2}\sum_{i=j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j},\qquad\hat{\nu}_{2}=\frac{6p}{2}-1,\qquad\hat{S}_{2}=\frac{1}{2}\sum_{i\neq j}\sum_{l=1}^{p}\widetilde{\bm{\Phi}}_{l,i,j}.$
-The last equality holds because the integrals in the third line are the kernels of inverse-gamma densities.
 Details about the Tuning Density Function
@@
 Figure C.1: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock.
-In this plot we estimate the shrinkage parameter $V_{b}$ using a hierachical Bayes approach.
-Figure C.2: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we use only data unitl 2019Q4.
+In this plot we estimate the shrinkage parameter $V_{b}$ using a hierarchical Bayes approach.
+Figure C.2: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we use only data until 2019Q4.
+i
 Figure C.3: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with eight lags instead of four.
@@
 Figure C.7: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock.
-In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (Bauer and Swanson, , 2023) instead of using the orthogonalized monetary policy surprise measure.
+In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (bauer2023reassessment) instead of using the orthogonalized monetary policy surprise measure.
 Figure C.8: The plot shows the posterior distributions of the impulse response of the trends to a monetary policy shock. In this plot we estimate the model with adding the excess bond premium.
 Figure C.9: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
-In this plot we estimate the shrinkage parameter $\lambda_{B}$ using a hierachical Bayes approach.
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
+In this plot we estimate the shrinkage parameter $\lambda_{B}$ using a hierarchical Bayes approach.
 Figure C.10: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
-In this plot we use only data unitl 2019Q4.
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
+In this plot we use only data until 2019Q4.
 Figure C.11: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
 In this plot we estimate the model with eight lags instead of four.
 Figure C.12: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
 In this plot we estimate the model with the one-year treasury yield instead of using the federal funds effective rate.
 Figure C.13: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
 In this plot we estimate the model with the two-year treasury yield instead of using the federal funds effective rate.
 Figure C.14: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
 In this plot we estimate the model with the personal consumption expenditure price index instead of the GDP deflator.
 Figure C.15: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
-In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (Bauer and Swanson, , 2023) instead of using the orthogonalized monetary policy surprise measure.
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
+In this plot we estimate the model with unadjusted monetary policy surprise (MPS) measure from (bauer2023reassessment) instead of using the orthogonalized monetary policy surprise measure.
 Figure C.16: The plot compares the actual trends and the counterfactual path with 68% credible bands.
-The counterfactual trends are calculated by shutting down all effect of all monetary policy shocks as is done in the calculation of historical decompositions, see Kilian and Lütkepohl, (2017).
+The counterfactual trends are calculated by shutting down the effects of all monetary policy shocks as is done in the calculation of historical decompositions, see kilian2017structural.
 In this plot we estimate the model with adding the excess bond premium.